### v1

# PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark

Abstract
While dense retrieval models have achieved remarkable success, rigorous evaluation of their sensitivity to the position of relevant information (i.e., position bias) remains largely unexplored. Existing benchmarks typically employ position-agnostic relevance labels, conflating the challenge of processing long contexts with the bias against specific evidence locations. To address this challenge, we introduce PosIR (Position-Aware Information Retrieval), a comprehensive benchmark designed to diagnose position bias in diverse retrieval scenarios. PosIR comprises 310 datasets spanning 10 languages and 31 domains, constructed through a rigorous pipeline that ties relevance to precise reference spans, enabling the strict disentanglement of document length from information position. Extensive experiments with 10 state-of-the-art embedding models reveal that: (1) Performance on PosIR in long-context settings correlates poorly with the MMTEB benchmark, exposing limitations in current short-text benchmarks; (2) Position bias is pervasive and intensifies with document length, with most models exhibiting primacy bias while certain models show unexpected recency bias; (3) Gradient-based saliency analysis further uncovers the distinct internal attention mechanisms driving these positional preferences. In summary, PosIR serves as a valuable diagnostic framework to foster the development of position-robust retrieval systems. [^1] https://github.com/Ziyang1060/PosIR
PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark
Ziyang Zeng1   Dun Zhang2   Yu Yan1 Xu Sun3Yudong Zhou2Yuqing Yang1thanks:   Corresponding author. 1Beijing University of Posts and Telecommunications 2Prior Shape 3Université Caen Normandie, ENSICAEN, CNRS, Normandie Univ, GREYC UMR6072, F-14000 Caen, France ziyang1060@bupt.edu.cn, dunnzhang0@gmail.com, yanyu2023@bupt.edu.cn xu.sun@unicaen.fr, zhouyudong@priorshape.com, yangyuqing@bupt.edu.cn
1 Introduction
The rapid advancement of dense retrieval has largely outpaced the evolution of rigorous evaluation protocols. While modern embedding models demonstrate impressive capabilities, characterizing their fine-grained behaviors requires benchmarks designed with higher diagnostic precision. In recent years, the Information Retrieval (IR) community has benefited from a series of milestone benchmarks. Early contributions like MS MARCO Nguyen et al. (2016) establish baselines for monolingual passage retrieval. Recognizing the linguistic diversity of the web, datasets such as MIRACL Zhang et al. (2023) extend the evaluation to multilingual scenarios. Additionally, benchmarks like BEIR Thakur et al. (2021), MMTEB Enevoldsen et al. (2025) and AIR-Bench Chen et al. (2025) aggregat diverse tasks to test general-purpose embeddings. These efforts have been instrumental in advancing the field.
Figure 1: The four-stage data generation pipeline of PosIR.
However, a critical dimension remains largely unexplored in existing benchmarks: position bias in information retrieval Hofstätter et al. (2021a). Specifically, current benchmarks generally operate on a position-agnostic assumption, providing relevance labels without specifying the location of the evidence. This conflation creates a diagnostic blind spot: when a retrieval model fails on a document, it is unclear whether the failure stems from the sheer volume of context (capacity limitation) or from bias against the location of the relevant span (positional sensitivity). Although Coelho et al. (2024); Zeng et al. (2025b) have observed primacy bias [^2] Primacy bias refers to models disproportionately focusing on the beginning of documents while overlooking relevant content appearing later. in embedding-based retrieval models, their experimental setups lack sufficient granularity to disentangle document length from information position. Without a controlled evaluation that orthogonalizes these variables, position bias cannot be strictly isolated from the model’s capacity to handle long contexts. Furthermore, the extent to which such bias varies across multilingual and cross-lingual retrieval settings remains an open question.
To address these gaps, we introduce PosIR (Position-Aware Information Retrieval), a comprehensive benchmark designed to rigorously evaluate and analyze position bias, which is characterized by three distinctive features: 1) Position-Aware: PosIR goes beyond coarse-grained document labels by associating each query with a precise reference span. By verifying relevance through strict span-based contrast, we enable the first systematic quantitative analysis of how the physical location of information impacts retrieval performance. 2) Heterogeneous: PosIR is designed to test robustness across diverse linguistic and semantic landscapes. It covers 31 domains and 10 languages (English, Chinese, and 8 translated languages), resulting in 310 datasets. This heterogeneity ensures that our findings on position bias are universal and not artifacts of a specific language or topic. 3) Length-Diverse: PosIR adopts a length-aware sampling strategy that spans documents from short passages to long texts. [^3] Up to 2048 tokens for native English and Chinese, with longer sequences for translated languages. This design enables the disentanglement of length and position factors, and exposes critical model behaviors, such as performance degradation in long-context settings, which are often overlooked by short-text benchmarks.
Using PosIR, we conduct extensive experiments on 10 popular IR models, yielding valuable insights into their performance characteristics:
  • Benchmarking Discrepancy: We find that retrieval performance on MMTEB correlates poorly with long-context retrieval performance on PosIR. Models that excel in short-context retrieval tasks often degrade significantly as document length increases.
  • Prevalence of Bias: Our evaluation reveals that position bias is pervasive and becomes more pronounced as document length increases, across both multilingual and cross-lingual retrieval settings. While most models exhibit primacy bias, favoring information at the beginning of documents, we identify unexpected recency bias in certain models.
  • Mechanistic Origins: Going beyond performance metrics, we employ gradient-based saliency analysis to uncover the internal mechanisms driving these biases. We identify two distinct attention behaviors that directly correlate with embedding models’ positional preferences over input tokens.
2 Benchmark Construction
The entire data generation pipeline of PosIR consists of four stages: 1) Bilingual corpora preparation, 2) Position-aware candidate generation, 3) Quality control, and 4) Multilingual translation. For detailed information of the models appearing in this paper, please refer to Table 12.
2.1 Bilingual Corpora Preparation
[…]
2.2 Position-Aware Candidate Generation
A retrieval dataset typically consists of three components: a document corpus, a set of queries, and relevance judgments (i.e., qrels). [^7] In TREC parlance, qrels refer to relevance judgments that indicate whether a document is relevant to a given query. After preparing the corpus in the initial stage, the candidate generation stage constructs the remaining two components of the dataset—queries and qrels—while explicitly accounting for the positional nature of relevance. As shown in Figure 1, the candidate generation process for a given domain proceeds as follows: 1) Sample one document from the corpus $\mathcal{D}$ as the positive document $d^{+}$. 2) Prompt LLMs to generate multiple queries $\{q_{i}\}_{i=1}^{m}$ for $d^{+}$ under a randomly sampled positional constraint (see Appendix C.1). 3) Prompt LLMs to locate the reference span $\mathrm{ref}_{i}$ corresponding to each generated question $q_{i}$ (see Appendix C.2). 4) Discard queries for which reference span localization fails (e.g., no unambiguous span can be identified or multiple candidate spans exist), and apply regular-expression matching to determine the character-level position span $\mathrm{pos}_{i}(\mathrm{start,end})$ of $\mathrm{ref}_{i}$ within $d^{+}$. In this stage, we use DeepSeek-V3.1 (Thinking Mode) DeepSeek-AI et al. (2025) for Steps 2 and 3, as these steps require multi-step reasoning over document content. After repeating the above steps thousands of times, we obtain the query set $\mathcal{Q}=\{q_{j}\}_{j=1}^{|\mathcal{Q}|}$, the positive document set $\mathcal{D}_{+}=\{d_{k}^{+}\}_{k=1}^{|\mathcal{D}_{+}|}$, the position-aware fine-grained qrels $\mathcal{R}_{+}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}|}$.
2.3 Quality Control
In this stage, we design a set of comprehensive quality control strategies to improve the reliability and precision of the generated dataset. As illustrated in Figure 1, the quality control process consists of two complementary components.
Reference Span Verification.
Since all reference spans in the qrels $\mathcal{R}_{+}$ are generated by LLMs, they may suffer from localization errors or hallucinations. To identify and remove such cases, we evaluate the necessity of each reference span using relevance contrast. Specifically, we employ two re-ranking models (e.g., bge-reranker-v2-minicpm-layerwise and Qwen3-Reranker-4B Zhang et al. (2025)) to compute two relevance scores for each query $q_{j}$: $s^{w/}_{j}$, the relevance score between $q_{j}$ and the full positive document $d_{j}^{+}$, and $s^{w/o}_{j}$, the relevance score between $q_{j}$ and a modified document obtained by removing the reference span $\mathrm{ref}_{j}$ from $d_{j}^{+}$. We retain only those queries for which $s^{w/}_{j}>s^{w/o}_{j}$ under both re-ranking models, indicating that the reference span is indispensable for establishing query-document relevance. Furthermore, for the remaining queries, we perform an additional verification step using an LLM-based relevance assessor (see Appendix C.3). Following the same contrastive setup, the LLM produces two 5-level relevance labels, $s^{\prime w/}_{j}$ and $s^{\prime w/o}_{j}$, ranging from 0 (completely irrelevant) to 4 (perfectly relevant). We retain a query only if $s^{\prime w/}_{j}>s^{\prime w/o}_{j}$ and $s^{\prime w/}_{j}\geq 3$, corresponding to a highly relevant judgment. Through this filtering process, we ensure that each query-document pair is relevant specifically due to a well-defined reference span at a particular position. This property is essential for accurately studying position bias in retrieval models.
Removing False Negatives.
[…]
After applying the above quality control process to each query, we obtain the refined query set $\mathcal{Q}^{\prime}$, corpus $\mathcal{D}^{\prime}$, and qrels $\mathcal{R}_{+}^{\prime}$, which together constitute the final bilingual dataset for each domain. Overall, this intentionally conservative quality control process prioritizes annotation precision over coverage, which is essential for isolating and analyzing fine-grained positional effects.
2.4 Multilingual Translation
After completing the earlier stages, we obtain high-quality bilingual English-Chinese datasets covering 31 domains, resulting in 62 position-aware retrieval datasets. To comprehensively study position bias in both multilingual and cross-lingual retrieval tasks, it is essential to consider a broad range of language scenarios. LLMs have exhibited remarkable proficiency in multilingual machine translation Zhu et al. (2024), significantly improving efficiency and eliminating the need to manually create position-aware retrieval datasets for each language from scratch. We employ Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a) to translate the English retrieval datasets across 31 domains into 8 languages, such as French, Spanish, Russian, and others, resulting in 248 datasets. Specifically, for each domain, we translate the English datasets $\mathcal{Q}_{en}^{\prime}$ and $\mathcal{D}_{en}^{\prime}$ into the target language datasets $\mathcal{Q}_{*}^{\prime}$ and $\mathcal{D}_{*}^{\prime}$, while the qrels $\mathcal{R}_{+}^{\prime}$ are shared across languages. This means that the position-aware relevance labels in the other languages remain consistent with the English dataset. To evaluate the translation quality, we employ both LLM-based automatic evaluation Kocmi and Federmann (2023) and human evaluation on a sampled subset, both of which demonstrate the overall high quality of the translations. However, during the multilingual translation process, we observe that token lengths grow disproportionately across languages due to richer morphology and longer sentence structures, which can cause the translated retrieval corpus $\mathcal{D_{*}}^{\prime}$ to exceed the predefined 2048-token length limit. Additionally, we observe issues such as invalid prefixes, refusal to translate, and repetitive or verbose translations. To further improve the translation quality, we implement a set of fixing and filtering strategies to ensure clarity and consistency in the results. For more details on the multilingual machine translation stage, please refer to Appendix B.
3 The PosIR Benchmark
Metric Value
Languages 10
Domains 31
Language-Domain Pairs 310
-    Avg #Queries 1,360
-    Avg #Documents 55,902
Total Queries Tokens 9,238,156
-    Avg Tokens per Query 21.91
Total Corpus Tokens 22,784,263,943
-    Avg Tokens per Document 1,314.75
Table 1: Summary of PosIR dataset statistics.
Table 1 provides a statistical overview of PosIR, which comprises 310 retrieval datasets spanning 10 languages [^8] Arabic, Chinese, German, English, French, Italian, Korean, Portuguese, Russian, Spanish. and 31 domains [^9] The 31 domains are grouped into the following high-level categories: Technology and Science, Industry and Economics, Health and Medicine, Society and Law, Entertainment and Culture, Finance, and Miscellaneous.. All datasets are following the same format as BEIR, i.e. corpus, queries and qrels, which are all available in the Hugging Face Hub. [^10] https://huggingface.co/datasets/infgrad/PosIR-Benchmark-v1 More details on PosIR are available in Appendix A.
3.1 Positional Distribution Analysis
In the position-aware fine-grained qrels $\mathcal{R}_{+}^{\prime}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}^{\prime}|}$, we explicitly incorporate positional information through $\mathrm{pos}_{j}$. To analyze the distributional characteristics of this positional information, we aggregate the qrels across all domains separately for the English and Chinese datasets. For each query-document pair, we compute the mean position of $\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})$ and normalize it by the character length of the corresponding positive document $d_{j}^{+}$, yielding a relative position in the range $[0,1]$. We then discretize the relative positions into 20 equal-width bins and count the number of queries falling into each bin. The resulting positional distributions are illustrated in Figure 2, confirming that reference spans are broadly distributed across document positions, without a pronounced bias toward early segments as observed in MS-MARCO Hofstätter et al. (2021b); Coelho et al. (2024). Such a broad distribution is crucial for evaluating retrieval models’ sensitivity to position bias, as it ensures that relevance signals occur throughout the document rather than being dominated by leading content.
3.2 Positioning of PosIR
We position PosIR as a multilingual, position-aware retrieval benchmark designed to expose position bias that is invisible to existing evaluation suites. Unlike popular retrieval benchmarks such as MTEB Muennighoff et al. (2023) and C-MTEB Xiao et al. (2024), which predominantly focus on short documents (typically under 512 tokens), PosIR explicitly incorporates documents with diverse and realistic length distributions, where relevant information may appear at arbitrary positions within the documents. A key distinguishing feature of PosIR is the fine-grained, position-aware relevance annotations. Rather than treating document relevance as a holistic property, PosIR associates each query-document pair with a localized reference span, allowing direct analysis of how retrieval models respond to the positional distribution of relevant information. This unique characteristic provides a solid empirical foundation for studying position bias in retrieval models, which is largely unexplored in existing benchmarks.
Figure 2: Distribution of normalized reference span positions in the English and Chinese datasets.
Model MMTEB PosIR Q1(512) Q2(1024) Q3(1536) Q4(2048) gte-multilingual-base Zhang et al. (2024) 57.16 47.37 61.28 48.79 39.21 32.01 bge-m3 Chen et al. (2024) 54.60 43.22 57.16 43.48 35.13 29.72 Qwen3-Embedding-0.6B Zhang et al. (2025) 64.65 53.63 62.93 54.41 48.11 43.32 inf-retriever-v1-1.5b Yang et al. (2025b) 62.96 58.81 67.82 58.40 53.90 51.20 Qwen3-Embedding-4B Zhang et al. (2025) 69.60 62.26 71.63 62.91 56.74 50.96 inf-retriever-v1 Yang et al. (2025b) 66.48 65.01 74.71 65.03 59.82 54.68 NV-Embed-v2 Lee et al. (2025) 56.72 45.02 70.48 47.12 26.33 16.27 llama-embed-nemotron-8b Babakhin et al. (2025) 68.69 64.09 75.76 64.16 57.47 52.28 Qwen3-Embedding-8B Zhang et al. (2025) 70.88 64.08 72.68 64.33 59.00 53.87 KaLM-Embedding-12B Zhao et al. (2025b) 75.66 51.87 74.01 54.64 35.11 27.65 Spearman Rank Correlation Coefficient (p-value) 0.62 (p=0.05) 0.73 (p=0.01) 0.71 ((p=0.02) 0.44 (p=0.2) 0.39 (p=0.2)
Table 2: nDCG@10 $\uparrow$ performance of 10 multilingual retrieval models on MMTEB (Retrieval Task) and PosIR. MMTEB scores are sourced from the official leaderboard. For PosIR, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). Spearman rank correlation coefficients between MMTEB and PosIR (overall and per length bucket) are reported at the bottom. “KaLM-Embedding-12B” denotes KaLM-Embedding-Gemma3-12B-2511. The detailed results for each language in PosIR are available in Table 10.
4 Experiments
In this section, we aim to address the following research questions:
RQ1: To what extent do model rankings on PosIR align with or diverge from those on established retrieval benchmarks, and in which scenarios does PosIR reveal discrepancies that are invisible under standard benchmarks?
RQ2: What systematic patterns of position bias (e.g., early vs. late relevance, document length effects, and cross-lingual variation) are exposed by the PosIR benchmark?
RQ3: What internal model behaviors and representational dynamics contribute to the emergence of position bias in neural retrieval models?
4.1 Comparison with MMTEB (RQ1)
Experimental Setup.
We adopt MMTEB (Retrieval Task) as a representative multilingual retrieval benchmark. We evaluate 10 popular IR models on both MMTEB and PosIR using nDCG@10, and compute the Spearman rank correlation coefficient between model rankings on the two benchmarks as a measure of consistency. To further analyze the effect of document length, we partition queries in PosIR into four length buckets (Q1–Q4) with 512-token intervals, based on the token length of their positive documents $d_{j}^{+}$ in the qrels $\mathcal{R}_{+}^{\prime}$. For translated datasets, token lengths are measured based on their original English counterparts to ensure consistent bucket assignment across languages. We then compute Spearman correlations between MMTEB and the average performance of models within each length bucket.
Main Results.
As shown in Table 2, the Spearman rank correlation between MMTEB and PosIR is 0.62 ($p=0.05$), indicating a moderate correlation. This suggests that while PosIR partially aligns with existing retrieval evaluations, it captures distinct performance characteristics not fully reflected by MMTEB. When controlling for document length, a clear trend emerges: the correlation is strongest for Q1 (documents up to 512 tokens), reaching 0.73 ($p=0.01$), and progressively decreases as document length increases. For Q4 (documents up to 2048 tokens), the correlation drops to 0.39 ($p=0.2$), indicating no statistically significant correlation. This length-dependent divergence implies MMTEB may primarily reflect model performance on short-document retrieval. In contrast, PosIR, through its length-aware sampling, exposes substantial discrepancies in how models handle long-context retrieval. Notably, most models exhibit pronounced performance degradation in Q4, despite nominally supporting long input contexts.
A distinct pattern is observed for NV-Embed-v2 and KaLM-Embedding-Gemma3-12B-2511, which achieve competitive performance on MMTEB and short-document queries (Q1), but suffer substantial degradation in Q3 and Q4. This behavior is consistent with their reported training configurations Lee et al. (2025); Zhao et al. (2025a), as both models are primarily trained with short-context inputs (e.g., up to 512 tokens), limiting their ability to effectively encode long documents. Furthermore, NV-Embed-v2, while strong in Q1, degrades severely in longer-document buckets in multilingual settings. This aligns with the fact that it is trained predominantly on English data, suggesting that long-context cross-lingual generalization poses a significant challenge when both document length and language shift are combined.
Model Dimension Attention Pooling PosIR Q1(512) Q2(1024) Q3(1536) Q4(2048) nDCG@10 $\uparrow$ PSI $\downarrow$ PSI $\downarrow$ PSI $\downarrow$ PSI $\downarrow$ Multilingual Retrieval gte-multilingual-base 768 bidirectional CLS 47.37 0.21 0.44 0.56 0.62 bge-m3 1024 bidirectional CLS 43.22 0.30 0.43 0.49 0.44 Qwen3-Embedding-0.6B 1024 causal last 53.63 0.21 0.38 0.47 0.49 inf-retriever-v1-1.5b 1536 bidirectional last 58.81 0.20 0.28 0.33 0.27 Qwen3-Embedding-4B 2560 causal last 62.26 0.13 0.30 0.39 0.44 inf-retriever-v1 3584 bidirectional last 65.01 0.14 0.21 0.17 0.18 NV-Embed-v2 4096 bidirectional mean* 45.02 0.19 0.53 0.73 0.81 llama-embed-nemotron-8b 4096 bidirectional mean 64.09 0.14 0.18 0.15 0.22 Qwen3-Embedding-8B 4096 causal last 64.08 0.12 0.23 0.31 0.34 KaLM-Embedding-12B 3840 bidirectional last 51.87 0.11 0.26 0.28 0.32 Average 55.54 0.18 0.32 0.39 0.41 Cross-lingual Retrieval gte-multilingual-base 768 bidirectional CLS 46.49 0.15 0.32 0.36 0.43 bge-m3 1024 bidirectional CLS 36.30 0.34 0.47 0.52 0.47 Qwen3-Embedding-0.6B 1024 causal last 49.66 0.14 0.37 0.48 0.51 inf-retriever-v1-1.5b 1536 bidirectional last 53.11 0.18 0.27 0.24 0.19 Qwen3-Embedding-4B 2560 causal last 60.76 0.13 0.30 0.43 0.47 inf-retriever-v1 3584 bidirectional last 62.87 0.13 0.22 0.19 0.16 NV-Embed-v2 4096 bidirectional mean* 45.97 0.16 0.14 0.52 0.62 llama-embed-nemotron-8b 4096 bidirectional mean 62.68 0.10 0.20 0.16 0.18 Qwen3-Embedding-8B 4096 causal last 61.93 0.12 0.26 0.41 0.45 KaLM-Embedding-12B 3840 bidirectional last 55.48 0.10 0.22 0.29 0.24 Average 53.53 0.16 0.27 0.36 0.37
Table 3: Position Sensitivity Index (PSI) $\downarrow$ of 10 multilingual retrieval models on PosIR. In both multilingual retrieval and cross-lingual retrieval (translated queries retrieving English documents) settings, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). “KaLM-Embedding-12B” denotes the “KaLM-Embedding-Gemma3-12B-2511” model, while NV-Embed-v2 improves the mean* pooling method with a latent attention layer. The detailed results for each language in PosIR are available in Table 10.
Figure 3: Mean nDCG@10 scores of 10 IR models across 20 relative position bins on the English subset of PosIR.
Figure 4: Mean nDCG@10 scores of 10 IR models across 20 relative position bins in the French-to-English cross-lingual setting of PosIR. “KaLM-Embedding-12B” denotes KaLM-Embedding-Gemma3-12B-2511.
4.2 Patterns of Position Bias (RQ2)
Experimental Setup.
Following Zeng et al. (2025b), we adopt the Position Sensitivity Index (PSI) as an intuitive diagnostic metric for quantifying position bias from a worst-case perspective. We divide the queries into four length buckets (Q1–Q4) according to the strategy in Section 4.1. Within each bucket, we calculate the average nDCG@10 scores across the 20 equal-width relative position bins (denoted as $\mathbf{s}=\{s_{1},\ldots,s_{20}\}$), following the discretization described in Section 3.1. Formally, PSI is defined as $\mathrm{PSI}=1-\frac{\min(\mathbf{s})}{\max(\mathbf{s})}$. A lower PSI indicates that the model’s performance is more consistent across document positions, reflecting reduced sensitivity to the location of relevant content. For a more detailed discussion of the PSI metric, please refer to the original work.
Main Results.
As shown in Table 3, we observe a clear trend: position bias tends to increase with document length across both multilingual and cross-lingual retrieval tasks. [^11] We consider a cross-lingual retrieval setting where non-English queries are used to retrieve an English corpus, which is a common scenario. For short documents (Q1), most models exhibit relatively low PSI values, suggesting that position bias is minimal when processing shorter inputs. However, as document length increases (Q3 and Q4), there is a marked rise in PSI for several models, indicating heightened sensitivity to position in long-context retrieval scenarios. Moreover, cross-lingual retrieval exhibits a lower overall PSI compared to the monolingual setting. We hypothesize this is partly because the generally lower performance in cross-lingual tasks compresses the score range, thereby reducing the discriminative power of the PSI metric across position bins. We also observe that for certain models, the PSI in Q3 or Q4 is counter-intuitively lower than in Q2. We attribute this to the models’ failure to effectively encode longer document representations, leading to uniformly degraded performance where position bias is no longer the dominant factor in score fluctuations. Moreover, we observe no clear correlation between positional sensitivity and architectural factors such as model size, vector dimension, attention mechanism (bidirectional / causal), or pooling strategy (CLS / last / mean). [^12] However, the latent attention layer preceding the mean pooling strategy may make NV-Embed-v2 more unique. This suggests that these architectural choices alone may not be the primary determinants of position bias, or their effects are overshadowed by training data distributions.
We visualize the fine-grained model performance across 20 relative position bins in Figures 3 and 4. Additional figures are available in the official GitHub repository. [^13] https://github.com/Ziyang1060/PosIR/tree/main/figs As shown in these figures, most models exhibit a generic pattern of primacy bias, where retrieval performance degrades as the relevant information moves towards the end of the document. An unexpected case is NV-Embed-v2, which exhibits a recency bias skewed towards the end of the document, diverging from the typical patterns observed in other models.
4.3 Mechanisms of Position Bias (RQ3)
Experimental Setup.
To better understand the mechanisms underlying the observed position bias, we conduct exploratory analyses on the internal behaviors of two representative models: Qwen3-Embedding-8B (exhibiting primacy bias) and NV-Embed-v2 (exhibiting recency bias). To quantitatively assess the contribution of tokens at different positions, we employ gradient-based saliency analysis Simonyan et al. (2014). Given a query $q$ and a relevant long document $d^{+}$ (with length $\geq$ 1024 tokens), where the reference span is located near the middle of the document (i.e., relative position within $[0.4,0.6]$), we compute the gradient of the cosine similarity score $s(q,d^{+})$ with respect to the input document embeddings. The magnitude of the gradient serves as a proxy for token importance, indicating how strongly each token influences the final document representation used for relevance matching. Specifically, for each document token $w_{i}$ at absolute position $i$, we compute the L2 norm of the gradient vector $\nabla_{w_{i}}s(q,d^{+})||_{2}$, which measures the sensitivity of the relevance score to that token. [^14] Special tokens at the sequence boundaries are excluded. To facilitate comparison across documents of varying lengths, we apply max normalization over all document tokens and rescale token positions to the range $[0,1]$ using linear interpolation, discretized into 100 relative position bins. We aggregate the saliency over 3,131 English query-document pairs that satisfy the length and position constraints to ensure statistical robustness.
Figure 5: Gradient-based saliency maps comparing the internal attention dynamics of Qwen3-Embedding-8B and NV-Embed-v2. The x-axis represents the normalized relative position within a document, and the y-axis shows the normalized L2 norm of the gradients. Shaded regions indicate one standard deviation.
Main Results.
We visualize the gradient-based saliency maps in Figure 5. The results reveal starkly contrasting internal behaviors between the two models. The saliency curve for Qwen3-Embedding-8B (Orange) is characterized by an extreme peak at the very beginning. This confirms that despite being a causal model, Qwen3 heavily relies on the initial tokens to establish the semantic context for the entire sequence. The subsequent sharp decay reflects the model’s inability to effectively propagate gradients from the middle sections, providing a mechanistic explanation for its tendency to ignore mid-document information. We hypothesize that the spike at the final position arises as a structural artifact of last-token pooling. In stark contrast, NV-Embed-v2 (Blue) exhibits a suppressed sensitivity to the beginning of the document, with gradients hovering near zero for the first 20% of tokens. Instead, it demonstrates a continuous rising trend starting from the mid-point (0.5) and peaking at the end. This J-shaped curve suggests that NV-Embed’s encoding mechanism prioritizes recent information, progressively overwriting or diluting early context. This explains its atypical recency bias and its failure to retrieve information located at the start of long documents.
5 Related Work
5.1 Position Bias in IR
Prior work has widely documented the existence of position bias in neural IR systems. Hofstätter et al. (2021b) identified that MS MARCO, a foundational IR dataset, is heavily skewed toward head-located relevant spans, leading fine-tuned models to inherit and amplify this position bias Jiang et al. (2021), which compromises the reliability of existing evaluations Rau et al. (2024). Recent studies confirm that such bias persists even in state-of-the-art embedding models Coelho et al. (2024); Fayyaz et al. (2025); Zeng et al. (2025b). Specifically, Zeng et al. (2025b) utilize answer start positions in SQuAD Rajpurkar et al. (2018) for bias analysis; however, SQuAD is constrained to English monolingual data and brief passages (avg. 117 words), hindering the generalization of findings to multilingual or long-context scenarios. Despite these insights, a standardized framework to rigorously evaluate position bias, especially within multilingual and long contexts, remains absent. Consequently, PosIR is proposed to bridge this gap.
5.2 Synthetic Data for IR
Leveraging LLMs for synthetic data generation has recently attracted growing attention in the IR community. Existing works can be broadly categorized into two primary use cases: training and evaluation. From a training perspective, LLMs are employed to address the scarcity of domain-specific or task-specific data. Methods like InPars Bonifacio et al. (2022) and Promptagator (Dai et al., 2023) prompt LLMs to generate synthetic queries for specific documents, facilitating effective retriever training in zero-shot settings. This paradigm has been further extended by Thakur et al. (2024); Wang et al. (2024) to support multilingual retrieval tasks. Regarding evaluation, recent studies have validated the reliability of synthetic test collections for benchmarking Rahmani et al. (2024). Khramtsova et al. (2024) leverage LLM-generated pseudo-labels and queries for unsupervised model selection on a target corpus. Furthermore, frameworks like AIR-Bench (Chen et al., 2025) provide automated pipelines to generate evaluation data for emerging domains efficiently. Inspired by these advancements, we extend the application of synthetic data generation to the specific challenge of position bias. Unlike prior works that focus on document-level relevance, we utilize LLMs to generate fine-grained, position-aware relevance annotations, which are crucial for isolating position bias. By incorporating rigorous quality control mechanisms, PosIR ensures the reliability of synthetic signals, enabling efficient and effective diagnosis of position bias in retrieval models.
6 Conclusion
In this paper, we introduce a new IR benchmark, PosIR, designed to diagnose position bias in information retrieval models. Unlike existing evaluation suites, PosIR leverages granular reference spans and length-diverse settings to effectively disentangle the effects of document length from the position of relevant information. Our experiments expose prevalent primacy and recency biases across both multilingual and cross-lingual retrieval settings, underscoring the limitations of current evaluation practices. By providing a reproducible pipeline and a diverse suite of 310 datasets, PosIR fills a crucial gap in the evaluation ecosystem, paving the way for more position-robust retrieval systems.
Limitations
Despite our efforts, this study has several limitations: 1) Although we employ LLMs to extract reference spans and implement a strict contrastive verification pipeline, a marginal risk remains that certain alternative reference spans might be overlooked. While our quality control is designed for high precision, such omissions could introduce noise into the position-aware evaluation. 2) While LLM-based translation significantly extends the linguistic diversity of PosIR, it is intrinsically difficult to completely eliminate subtle semantic drift or translation artifacts across all datasets. Although our sampling-based automatic and human evaluations indicate high fidelity, translation quality in certain low-resource languages might still be influenced by the underlying model bias. 3) Our empirical analysis focuses on embedding-based dense retrieval models. Although prior research Zeng et al. (2025b) indicates that cross-encoders may exhibit superior position robustness, a systematic large-scale investigation of alternative architectures, including generative retrieval, remains for future work. Nevertheless, PosIR provides the necessary model-agnostic framework to facilitate such benchmarking. 4) Our investigation is currently limited to text retrieval. Whether and how position bias manifests in multimodal retrieval contexts (e.g., text-to-image or video retrieval) remains an open question. We leave the exploration of these cross-modal positional dynamics to future work.
Ethical Considerations
The construction of PosIR involves text generation by LLMs. Consequently, the synthetic components of the datasets (e.g., queries) may inherently reflect the socio-cultural biases, stereotypes, or toxicity present in the base models. Additionally, as the document corpora are derived from real-world open-source datasets (IndustryCorpus2 and FineWeb), they may inevitably contain personally identifiable information or sensitive content. To mitigate these risks, we emphasize that PosIR is intended solely for research benchmarking and evaluation purposes. We strictly prohibit its use for training generative models to prevent the propagation of potential toxicity or hallucinations.
Appendix A PosIR Datasets
A.1 Overview
PosIR comprises 310 datasets (17,329,673 documents and 421,708 queries) spanning 10 languages and 31 domains, forming a large-scale multilingual benchmark for information retrieval. For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels, which are all available in the Hugging Face Hub. [^15] https://huggingface.co/datasets/infgrad/PosIR-Benchmark-v1 Table 4 summarizes language-level statistics, showing well-balanced corpus sizes across languages, with each language contributing approximately 1.7M–1.9M documents. Token length statistics reveal substantial cross-lingual variation that reflects differences in writing systems and morphological complexity. Logographic languages exhibit more compact representations (e.g., Chinese and English), whereas morphologically rich alphabetic languages require longer sequences (e.g., Arabic and Russian). Table 5 reports domain-level aggregates and highlights pronounced variation in domain scale and document characteristics. Domain sizes range from several hundred thousand documents to over three quarters of a million, while token length distributions reflect domain-specific content properties. Technical domains such as Information Security tend to contain longer documents, whereas more formal or symbolic domains, such as Mathematics and Statistics, exhibit comparatively shorter texts. This diversity enables evaluation of IR systems under heterogeneous conditions. For comprehensive coverage, Tables 1343 provide detailed statistics for all 31 domains, presented individually for each domain. These tables report language-specific document counts, query counts, average token lengths, and token length ratios relative to English. Such fine-grained reporting supports reproducibility and facilitates the analysis of domain–language interaction effects in multilingual IR.
Document Count Token Length
Domain Corpus Queries Ratio Min Q1 Median Q3 Max Mean
Accommodation Catering Hotel 439,993 14,612 2.56% 10 607 1173 1886 4432 1294.6
Aerospace 581,933 13,546 3.36% 9 610 1207 1955 4585 1326.1
Agriculture Forestry Animal Husbandry Fishery 581,329 16,175 3.37% 9 589 1231 2013 4719 1346.3
Artificial Intelligence Machine Learning 551,318 11,154 3.17% 5 538 1113 1884 4540 1276.7
Automobile 565,317 13,047 3.26% 5 584 1120 1870 4346 1271.3
Biomedicine 616,910 13,948 3.55% 7 580 1195 1972 4709 1327.3
Computer Communication 620,254 14,144 3.57% 6 607 1172 1908 4544 1300.5
Computer Programming Code 519,815 6,522 2.97% 3 701 1201 1853 4127 1307.0
Current Affairs Government Administration 606,998 12,832 3.49% 2 619 1225 1985 4785 1341.8
Electric Power Energy 599,389 13,119 3.45% 6 613 1220 2001 4676 1352.1
Film Entertainment 596,575 16,911 3.46% 10 592 1195 1906 4605 1295.0
Finance Economics 607,364 12,948 3.49% 5 626 1227 1982 4571 1343.5
Fineweb 698,633 16,748 4.03% 5 575 1175 1919 4645 1301.7
Fire Safety Food Safety 313,846 12,852 1.84% 13 672 1197 1953 4612 1361.1
Game 442,997 13,014 2.57% 6 646 1234 1970 4634 1339.5
Law Judiciary 603,544 12,947 3.47% 10 626 1221 1971 4676 1338.7
Literature Emotion 598,538 15,717 3.46% 3 613 1181 1915 4653 1299.6
Mathematics Statistics 618,864 10,324 3.54% 2 588 1140 1827 4180 1244.3
Medicine Health Psychology Traditional Chinese Medicine 633,903 13,925 3.65% 14 609 1211 1987 4828 1342.9
Mining 485,528 13,345 2.81% 9 510 1114 1890 4771 1266.1
News Media 539,996 12,879 3.12% 5 588 1230 1964 4738 1326.3
Other Information Services Information Security 321,187 8,590 1.86% 13 669 1246 2026 4402 1384.2
Other Manufacturing 596,244 16,025 3.45% 6 594 1131 1884 $44$ 1283.5
Petrochemical 577,348 12,803 3.32% 6 570 1180 1962 4583 1316.6
Real Estate Construction 567,575 16,259 3.29% 13 594 1211 1963 4693 1328.5
Sports 566,912 13,787 3.27% 13 565 1163 1892 4531 1281.9
Subject Education Education 618,409 16,796 3.58% 9 625 1221 1975 4569 1339.0
Technology Scientific Research 614,697 14,198 3.54% 3 595 1195 1960 4592 1318.4
Tourism Geography 549,553 14,974 3.18% 5 560 1176 1927 4633 1295.9
Transportation 547,312 14,129 3.16% 10 573 1190 1965 4640 1323.3
Water Resources Ocean 547,392 13,438 3.16% 7 571 1194 1960 4563 1325.8
Total/Avg 17,329,673 421,708 100.00% 7 600 1189 1939 4581 1316.1
Table 5: Summary of corpus statistics aggregated by domain.
A.2 Diversity Analysis
Query Diversity
To analyze the diversity of query types in PosIR, we adopt a keyword-based heuristic that categorizes queries according to their leading interrogative terms (e.g., what), with non-interrogative queries grouped as claim. Table 6 summarizes the distribution of query types in the English and Chinese retrieval datasets. [^16] The OTHERS category mainly includes yes/no questions—typically those starting with auxiliary verbs (e.g., do/does/did, can, will) or be verbs (e.g., is/are)—as well as other infrequent or implicit question forms that do not match predefined interrogative patterns. Several observations can be made from the results. First, what queries constitute the largest proportion of the datasets, followed by how and why queries, indicating that the generated queries predominantly seek factual explanations and procedural information. Notably, Chinese queries contain a higher proportion of why and who, whereas English queries exhibit a larger share of queries categorized as others, suggesting differences in query formulation across languages. In addition, a small fraction of queries are categorized as claim, representing declarative statements that express information needs without explicit interrogative forms.
Figure 6: Inter-domain similarity heatmap. The lower triangular part of the matrix shows similarity scores computed from the English corpus, while the upper triangular part corresponds to those from the Chinese corpus. The diagonal is omitted for clarity.
Corpus Diversity
Following the approach in MTEB Muennighoff et al. (2023), we compute inter-domain similarities separately for English and Chinese corpora, each consisting of 31 domains, in PosIR. Specifically, we adopt the highest-scoring model on the MTEB (eng, v2) STS task, QZhou-Embedding (91.65) Yu et al. (2025), to generate embeddings for all documents within each domain. We then average document embeddings at the domain level and compute pairwise cosine similarities between domains. The resulting similarity matrix is visualized as a heatmap in Figure 6.
Query Type English (%) Chinese (%)
WHAT 34.46 41.44
HOW 26.60 17.93
WHY 12.33 23.18
WHICH 5.53 1.75
WHEN 3.14 5.68
WHERE 2.96 1.86
WHO 2.67 5.25
CLAIM 0.42 0.03
OTHERS 11.88 2.90
Total 100.00 100.00
Table 6: Distribution of query types in the English and Chinese retrieval datasets.
Appendix B Multilingual Machine Translation
In this section, we detail the selection process of the translation model and the rigorous quality control protocols applied to construct the multilingual subset of PosIR.
Data Construction.
To construct a representative evaluation set for translation quality, we perform stratified sampling from the 31 domain-specific English corpora utilized in PosIR. Within each domain, we select 12 documents stratified by token length into four bins (0–512, 512–1024, 1024–1536, and 1536–2048 tokens, measured by the Qwen3 tokenizer), yielding a total of 372 source texts. This design ensures balanced domain coverage while capturing a broad spectrum of document lengths and linguistic complexities.
Translation Models.
We evaluate four representative translation systems: (1) Google Translate [^17] https://translate.google.com, a widely used commercial neural machine translation service; (2) Hunyuan-MT-7B Zheng et al. (2025), a specialized translation model achieved first place in the WMT25 competition; (3) Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a), a 31B-parameter open-source general MoE model; and (4) GPT-4o OpenAI et al. (2024), a frontier general model serving as a strong baseline. Each system translates the sampled texts into 9 target languages: Arabic (ara-Arab), French (fra-Latn), German (deu-Latn), Italian (ita-Latn), Korean (kor-Kore), Portuguese (por-Latn), Russian (rus-Cyrl), and Spanish (spa-Latn). The prompt for the translation task is provided in Appendix C.4.
B.1 Automatic Evaluation
B.1.1 Evaluation Protocol
To efficiently evaluate the translation quality across 8 languages and 31 domains, we employ an LLM-based reference-free evaluation approach Kocmi and Federmann (2023) (see Appendix C.5). Specifically, we utilize GPT-4o as an automatic evaluator to assess translation quality along four dimensions: Fluency (naturalness and readability of the target text), Accuracy (semantic fidelity to the source), Completeness (preservation of all source information), and Style Consistency (appropriateness of tone and register). Each dimension is rated on a 5-point Likert scale (1 = Poor, 5 = Excellent), and an overall score is computed as the mean of the four dimensions. In total, this results in 2,976 evaluated translation outputs (372 samples $\times$ 8 languages).
Language Qwen3-30B-A3B Google Translate Hunyuan-MT-7B GPT-4o
Flu Acc Com Sty Ove Flu Acc Com Sty Ove Flu Acc Com Sty Ove Flu Acc Com Sty Ove
Arabic 4.38 4.18 4.19 4.33 4.27 4.33 4.24 4.26 4.33 4.29 4.30 3.70 3.28 4.14 3.85 4.25 4.16 4.14 4.22 4.19
French 4.85 4.63 4.54 4.84 4.71 4.64 4.47 4.43 4.64 4.54 4.40 3.74 3.27 4.24 3.91 4.57 4.48 4.38 4.56 4.50
German 4.61 4.45 4.51 4.60 4.54 4.58 4.35 4.38 4.55 4.47 4.40 3.74 3.33 4.21 3.92 4.55 4.50 4.45 4.55 4.51
Italian 4.67 4.48 4.44 4.65 4.56 4.39 4.37 4.37 4.39 4.38 4.37 3.67 3.15 4.14 3.83 4.44 4.34 4.32 4.40 4.37
Korean 4.39 4.12 4.13 4.35 4.25 4.24 4.17 4.25 4.24 4.23 4.40 3.91 3.54 4.28 4.03 3.88 3.74 3.78 3.87 3.82
Portuguese 4.81 4.61 4.59 4.81 4.71 4.59 4.45 4.47 4.56 4.52 4.42 3.75 3.29 4.24 3.93 4.71 4.60 4.53 4.70 4.63
Russian 4.69 4.44 4.47 4.68 4.57 4.51 4.28 4.30 4.45 4.39 4.30 3.57 3.14 4.11 3.78 4.35 4.23 4.19 4.33 4.28
Spanish 4.77 4.60 4.61 4.77 4.69 4.52 4.39 4.40 4.52 4.46 4.37 3.71 3.22 4.22 3.88 4.63 4.52 4.49 4.62 4.57
Average 4.65 4.44 4.43 4.63 4.54 4.47 4.34 4.36 4.46 4.41 4.37 3.72 3.28 4.20 3.89 4.42 4.32 4.29 4.41 4.36
Table 7: Automatic evaluation of multilingual translation quality by languages and dimensions (1-5 scale). Evaluation dimensions include Fluency (Flu), Accuracy (Acc), Completeness (Com), Style Consistency (Sty), and Overall Average Score (Ove). “Qwen3-30B-A3B” denotes the “Qwen3-30B-A3B-Instruct-2507” model.
Figure 7: Cross-lingual translation quality heatmap (using Overall Average Score) across 3 translation models and 5 language families. “Qwen3-30B-A3B” denotes the “Qwen3-30B-A3B-Instruct-2507” model.
B.1.2 Evaluation Results
Table 7 present aggregated evaluation results across four translation models and five quality dimensions. Qwen3-30B-A3B-Instruct-2507 achieves the highest overall score (4.54), followed by Google Translate (4.41), GPT-4o (4.36), and Hunyuan-MT-7B (3.89). Dimension-wise analysis reveals that Qwen3-30B-A3B-Instruct-2507 ranks first across all metrics, demonstrating particular strengths in fluency (4.65) and style consistency (4.63), while maintaining competitive performance in accuracy (4.44) and completeness (4.43). Google Translate exhibits a balanced profile with relatively uniform scores across dimensions, indicating stable but less exceptional translation quality. In stark contrast, Hunyuan-MT-7B shows severe degradation in completeness (3.28) and accuracy (3.72), with performance gaps exceeding 1.0 point compared to top-performing models. Despite achieving comparable fluency score (4.37), qualitative analysis reveals that Hunyuan-MT-7B frequently sacrifices semantic fidelity to preserve surface-level linguistic smoothness, resulting in content omissions and meaning distortions, particularly when handling complex or information-dense source texts. Moreover, as shown in Figure 7, a cross-linguistic analysis reveals that Romance languages such as French achieve relatively higher translation quality, whereas Arabic and Korean exhibit comparatively lower performance.
B.2 Human Evaluation
Initial inspection using automatic evaluation suggests that Qwen3-30B-A3B-Instruct-2507 produces high-quality translations across PosIR, achieving competitive performance in comparison to other methods. However, given that translation quality appears strong, rigorous human validation is essential. While automatic evaluation can detect catastrophic failures, it may be insensitive to subtle but critical errors in high-quality translations, such as nuanced mistranslations, contextual inappropriateness, or domain-specific inaccuracies (Freitag et al., 2021). Due to limited resources, we conduct focused human evaluation on three typologically diverse languages (Arabic, French, and Russian) to provide expert assessment of translation quality, facilitating a direct comparison between human evaluation and LLM-based assessments.
Dimension French Russian Arabic Avg.
Fluency 4.71 4.65 4.58 4.65
Accuracy 4.52 4.48 4.35 4.45
Completeness 4.89 4.82 4.76 4.82
Style Consistency 4.68 4.61 4.52 4.60
Overall Avg. Score 4.70 4.64 4.55 4.63
Table 8: Human evaluation scores by languages and dimensions (1-5 scale).
B.2.1 Annotator Recruitment
Five annotators were recruited from graduate programs in linguistics, natural language processing, and applied foreign languages via established academic networks. All annotators were advanced graduate students or recent PhD graduates with expertise in computational linguistics and/or translation studies. For Arabic and Russian, native speakers of the respective languages conducted the annotations. For French, two annotators holding DALF C1/C2 certificates (Diplôme Approfondi de Langue Française, the highest level of French language certification) completed the evaluation. All annotators were compensated at the French SMIC (Salaire Minimum Interprofessionnel de Croissance) hourly rate of €12.02 (\$14.15 USD as of December 2025), in compliance with French labor regulations.
B.2.2 Evaluation Protocol
Figure 8: Human vs. GPT-4o score comparison by language. Evaluation dimensions include Fluency (Flu), Accuracy (Acc), Completeness, and Style Consistency. The overlapping radar shapes indicate strong alignment in quality assessment patterns.
Figure 9: Confusion matrix of human vs. GPT-4o scores. Each row shows distribution of GPT-4o scores for a given human score. Data concentrates in the high-score diagonal (4–5), with GPT-4o showing systematic strictness for top-rated samples.
Annotators evaluated 372 translation pairs across 3 languages (1,116 samples in total) generated by Qwen3-30B-A3B-Instruct-2507. The evaluations were conducted in accordance with the consistency standards outlined in the automatic evaluation, as described in Appendix B.1.1. Annotators were given both the English source text and the target language translation, with no access to reference translations to prevent anchoring bias.
B.2.3 Evaluation Results
Table 8 presents the mean scores across languages and dimensions. All language-dimension combinations achieved scores between 4.35 and 4.89, indicating consistently high translation quality. Over 97% of translations received scores $\geq$4 across all dimensions. Several patterns emerge from the results. First, completeness scores consistently achieve the highest ratings (4.76–4.89), indicating that Qwen3-30B-A3B-Instruct-2507 rarely omits source information and preserves semantic integrity well. Second, fluency scores consistently exceed accuracy scores across all languages (4.58–4.71 vs. 4.35–4.52), suggesting the model generates natural-sounding translations in target languages while maintaining acceptable fidelity. Third, language-specific differences reflect the typological distance from English: French achieves the highest average scores (4.70), likely due to abundant training data and linguistic proximity; Arabic shows relatively lower accuracy (4.35), potentially due to its complex morphological system; Russian performs in between (4.64), despite its challenging case system. The score distribution exhibits a pronounced ceiling effect: 67.1% of samples received 5 (Excellent), 30.4% received 4 (Good), and only 2.5% scored $\leq$3. This confirms that the Qwen model produces high-quality translations suitable for PosIR, with >>97% of samples rated Good or Excellent.
Dimension Agreement (±1) MAE
Fluency 98.1% 0.44
Accuracy 96.9% 0.51
Completeness 90.6% 0.57
Style Consistency 96.9% 0.39
Table 9: Human-GPT-4o agreement within ±1 point tolerance.
B.3 GPT-4o Human Evaluation Alignment
We validate GPT-4o as a translation evaluator by comparing its ratings against human scores on the same 1,116 samples. Figure 8 shows the score distribution comparison between human annotators and GPT-4o across languages. The radar charts demonstrate high shape alignment: both raters identify completeness as the strongest dimension and accuracy as relatively lower, with mean differences within 0.3 points across all dimensions. Under a ±1 tolerance threshold (appropriate for ordinal Likert scales), GPT-4o achieves 96.15% agreement with human ratings. Table 9 shows per-dimension results. Figure 9 presents the detailed discrepancy analysis, revealing two notable patterns. In the high-score region, data concentrates along the 4–5 diagonal, explaining the strong agreement rates. However, for samples rated 5 by human annotators, GPT-4o averages 4.5–4.7, exhibiting a systematic offset of 0.3-0.3 to 0.5-0.5 points. This suggests GPT-4o acts as a stricter judge, more inclined to identify potential flaws than to endorse perfection, effectively serving as a conservative lower-bound estimator. Conversely, in the sparsely populated low-score region, GPT-4o tends toward leniency: for the rare samples rated 1–3 by humans (only 2.5% of data), GPT-4o often assigns 4–5 scores, suggesting potential blind spots in detecting severe semantic errors. Fortunately, such low-quality samples are rare in our dataset, so GPT-4o’s conservative behavior on high-quality translations, where the vast majority of data resides, provides reliable quality assurance for PosIR. In summary, the strong human-GPT-4o alignment validates automated quality verification for the full dataset, establishing confidence in the translation reliability of the Qwen model.
B.4 Model Selection Rationale
Based on comprehensive evaluation results and practical considerations, we select Qwen3-30B-A3B-Instruct-2507 as the primary translation model for constructing the PosIR multilingual corpus. This decision is justified by three key factors: (1) superior translation quality, consistently achieving the highest scores across all evaluation dimensions and target languages, with an overall advantage of 0.13 points over Google Translate and 0.21 points over GPT-4o; (2) strong cross-lingual robustness, maintaining stable performance across diverse linguistic typologies and writing systems without significant quality degradation; (3) favorable resource efficiency, delivering high translation quality with moderate computational requirements, which enables large-scale multilingual data construction under practical inference-time constraints. Additionally, as an open-source and freely accessible model, it offers further flexibility and cost-effectiveness for large-scale deployment and application.
B.5 Fixing and Filtering Strategies
Through manual inspection of the translated outputs, we observed recurring artifacts that could degrade data quality, such as invalid prefixes, refusal-style responses, and repetitive or verbose translations. To systematically mitigate these issues while preserving corpus-query-qrels consistency, we design and apply a three-stage cleaning pipeline.
B.5.1 Prefix Removal
The Qwen3-30B-A3B-Instruct-2507 model sometimes prepends explanatory text (e.g., “Here is the translation:”) to its outputs. To identify and remove such artifacts, we adopt a dual-strategy prefix detection heuristic. For corpus texts, we employ a co-occurrence pattern matching strategy that identifies the presence of translation-related keywords (“translate” or “translation”) and newline characters within a proximity window of 500 characters, starting from the first 500 characters and extending up to 2,000 characters. For queries, which are typically shorter, we apply full-text matching. Candidate prefixes are validated as English using langdetect [^18] https://github.com/fedelopez77/langdetect and manually reviewed before compilation into a blacklist (118 patterns). Overall, this procedure affected 54 instances across 40 files, accounting for 0.0003% of the 17.7M records and spanning 8 non-English languages.
B.5.2 Refusal Detection
We employ a two-stage filtering approach to identify translation refusals. In the first stage, we apply lightweight blacklist matching combined with language detection to flag candidate refusals, prioritizing high recall with minimal computational overhead. In the second stage, flagged instances are validated using Qwen3-Max [^19] https://qwen.ai/blog?id=qwen3-max, which inspects the first 1,500 characters to distinguish refusal-related meta-commentary (e.g., “I cannot translate”) from legitimate translations. This process identified 2,249 refusals, predominantly in Arabic (1,378 instances) and Korean (323). We remove 2,223 corpus entries and 26 queries, with cascading deletion of 26 corresponding qrels to maintain data consistency.
B.5.3 Statistical Outlier Filtering
We perform token-length-based outlier filtering at the language–domain level using the Qwen3 tokenizer. For each stratum, we compute the interquartile range (IQR) and define asymmetric thresholds, with a lower bound of $Q1-3.0\times IQR$ and an upper bound of $Q3+1.5\times IQR$. This asymmetric design focuses on filtering abnormally long texts, which are more likely to result from translation errors, duplication, or unintended document concatenation. It enforces a stricter upper bound to remove such cases, while allowing reasonable length variation among shorter documents. We process only corpus outliers, cascading deletions to associated queries and qrels. In total, 15,641 corpus documents (0.11%) and 332 queries (0.10%) are removed. The largest reductions occur in Arabic (5,490 corpus entries; 0.25%) and Korean (4,082; 0.24%), followed by Russian (1,557; 0.11%). Overall, the removal ratios remain consistently low across all languages, indicating that the filtering procedure effectively eliminates anomalous long texts while preserving the structural integrity and cross-lingual balance of the benchmark.
Appendix C Prompts
C.1 Prompt for Position-Aware QA Generation
We use the following prompt to instruct an LLM to generate diverse and realistic question–answer (QA) pairs under specific positional constraints. Unlike explicitly segmenting the document into segments, we always feed the entire document to the LLM. Position control is achieved solely by injecting a sampled positional constraint into a dedicated slot in the prompt.
Sampling and Constraint Injection.
For each generation instance, we randomly sample one positional constraint from a predefined set and insert it into the prompt via Positional Requirement Slot. This constraint guides the LLM to produce questions whose answers primarily come from a particular region of the document, while the model still has access to the complete document context.
The Positional Requirement Slot is filled with exactly one of the following options:
  • Focus on first third: The question should primarily reference the first third of the document.
  • Focus on middle third: The question should primarily reference the middle third of the document.
  • Focus on final third: The question should primarily reference the final third of the document.
Prompt Usage.
The prompt below enforces key requirements including answerability, naturalness, and the avoidance of meta references to the source document. The positional constraint is an internal control signal and must not be mentioned explicitly in the generated questions. The question type list is intentionally over-complete to support diverse generation; not all types are expected to be instantiated for every document.
# Task Description Generate user questions in realistic search engine scenarios based on a given document. # Core Requirements 1. Answerability: Answers must be explicitly stated or directly inferable from the document (no outside knowledge). 2. Authenticity: Questions must reflect natural human phrasing. 3. No Referencing: Avoid terms like “this text”, “the author”, or “this document”. 4. {{Positional Requirement Slot}} # Question Configuration Configure two elements: Question Type and Expression Style. Select configurations contextually. ## Question Type 1. **Fact Query**: Retrieve specific information (e.g., dates, metrics, definitions). 2. **Operation Guide**: Request steps/methods to accomplish tasks (focuses on actionable steps). 3. **Cause Analysis**: Investigate reasons behind occurrences or phenomena. 4. **Comparison Choice**: Contrast differences/advantages to facilitate decisions. 5. **Concept Explanation**: Clarify meanings, categories, or structures of terms/concepts. 6. **Viewpoint Verification**: Validate accuracy/reliability of claims/data (requires credibility evaluation). 7. **Prediction/Speculation**: Inquire about future outcomes or plausible inferences strictly grounded in the document content. 8. **True/False Judgment**: Questions answerable with "yes" or "no". 9. **Background Exploration**: Understand context, history, current status, or related environments. 10. **Seeking Advice**: Request subjective opinions, experiences, or recommendations (emphasizes actionable suggestions). 11. **Open Discussion**: Pose broad/debatable topics to stimulate dialogue. 12. **Emotional Resonance**: Express feelings (frustration/seeking help/excitement) to obtain empathy/support. 13. **Feasibility Assessment**: Query viability, costs, or resource requirements for plans/ideas. 14. **Mathematical Reasoning**: Problems needing numerical calculations or logical deductions (require step-by-step solutions). 15. **Content Creation**: Request original text, narratives, or creative output. 16. **Relationship Counseling**: Focuses on conflict resolution in interpersonal interactions, communication strategies, or emotional advice. 17. **Diagnostic Troubleshooting**: Analyzes the root causes of issues based on anomalies and provides solutions. 18. **Hypothetical Scenarios**: Explores possibilities, ethical implications, or logical inferences under fictional or extreme conditions. 19. **Decision Advisory**: Assists in analyzing pros and cons, weighing factors for significant or personalized choices. 20. **Personal Growth**: Requests specific strategies, methods, or learning path planning for skill development, habit formation, efficiency enhancement, goal setting, or mindset adjustment. ## Expression Style - Concise | Casual | Informal - Formal | Professional - Technical | Academic # Steps Follow the steps below and output results for each step in order. 1. Scenario Setup: Envision potential users and contexts where this document would be relevant. 2. Type Selection: Identify appropriate question types based on the content emphasized by the sampled positional constraint. 3. Style Matching: Select compatible expression styles. 4. Final Generation: Generate the final QA pairs and present them as a JSON array enclosed in a Markdown code block.
    "question": "Generated question",
    "answer": "Answer text"
# Important Tips 1. Users do not know the document content. Questions must (a) strictly avoid document references, and (b) use natural phrasing. 2. The positional constraint is for generation control only; do not reveal or mention it in the questions. 3. For subjective question types (e.g., advice), answers must be grounded in the document’s stated information; do not introduce external opinions. 4. If the document quality prevents question generation, output: “Document quality insufficient for generation.” Now, let’s start this task: <document>{{DOCUMENT}}</document>
C.2 Prompt for Reference Locating
Although explicit positional constraints are imposed during QA generation, the resulting questions may still violate these constraints in practice. To address this issue, we design the following prompt to instruct an LLM to locate supporting reference text within the original document, thereby enhancing the factual grounding and answerability of the generated questions.
# Task Description Given a list of questions and a document, for each question, **precisely identify and extract continuous text segments** from the document that **directly answer or support answering** this question. ## Requirements 1. **Text Integrity**: Copy document content verbatim. Do not paraphrase, summarize, or splice across paragraphs. Preserve original punctuation, whitespace, and line breaks exactly. 2. **Extraction Principles**: Prioritize complete clauses/semantic units as minimal extraction units; If multiple relevant segments exist, output all segments completely without omission. 3. **Relevance Criteria**: Extracted segments must satisfy **at least one** of: (a) Containing direct answers to the question; (b) Providing essential contextual support for answering the question 4. **Question Coverage**: Process **every question** in the provided question list. 5. **Accurately Assess Question Answerability**: If the document can answer the question (directly or indirectly), extract the reference text as supporting evidence; If the document cannot provide useful information, return an empty list. ## Output Specification: Use strictly this JSON format.
    "question": "Original Question",
    "reference_texts": [
      "Document Segment 1",
      "Document Segment 2",
Now, let’s start this task: <document>{{DOCUMENT}}</document> <questions>{{QUESTIONS}}</questions>
C.3 Prompt for Relevance Assessment
Thomas et al. (2024) demonstrate that large language models such as GPT-4 can achieve labeling quality comparable to human annotators when producing gold-standard relevance labels for search systems. Moreover, Zhuang et al. (2024); Zeng et al. (2025a) show that incorporating fine-grained relevance labels into prompts for LLM-based rerankers significantly improves zero-shot reranking performance by better distinguishing borderline and ambiguous relevance cases. Motivated by these findings, we employ DeepSeek-V3.1 (Non-thinking Mode) [^20] https://api-docs.deepseek.com/news/news250821 as a relevance labeler, adopting a 5-level relevance labeling scheme to capture fine-grained distinctions in relevance. This labeling process is used solely for auxiliary verification rather than primary relevance annotation.
# Task Given a list of questions and a document, evaluate the relevance between each question and the document on a 0-4 scale: 0 = Completely irrelevant 1 = Slightly relevant (minimal connection) 2 = Moderately relevant (The document addresses parts of the question but lacks completeness, depth, or direct alignment) 3 = Highly relevant (covers most aspects of the question) 4 = Perfectly relevant (can correctly address the question) # Output Format Output all the relevance scores in the **same order** as the input questions in the format of a JSON array in Markdown. An example, assuming there are three questions, then your output is:
  1,
  3,
  0
Now, let’s start this task: <document>{{DOCUMENT}}</document> <questions>{{QUESTIONS}}</questions>
C.4 Prompt for Multilingual Machine Translation
You are a world-class professional translator, fluent in both English and {{LANGUAGE}}. Your task is to translate the following text into {{LANGUAGE}}. Please adhere to the highest standards of translation, ensuring the output is: 1. **Faithful:** Accurately convey the original meaning, content, and intent without omission or distortion. 2. **Expressive:** Make the translation smooth, natural, and easy to understand for a native speaker of the {{LANGUAGE}}. 3. **Elegant:** Preserve the style, tone, and literary grace of the original text, making it aesthetically pleasing to read. Here is the text: {{TEXT}}
C.5 Prompt for Translation Quality Evaluation
# Task You are a professional translation quality evaluator. Evaluate the English to {{LANGUAGE}} translation using a 5-point integer scale (1-5, whole numbers only). ## Evaluation Criteria 1. **Accuracy**: Semantic fidelity to original meaning - no mistranslations, omissions, or additions. 2. **Fluency**: Natural language quality, grammar, and readability in target language. 3. **Completeness**: All source content translated without missing information or concepts. 4. **Style Consistency**: Appropriate tone, register, and style matching the source text context. ## Scoring Scale - **5**: Excellent (professional publication quality, no issues). - **4**: Good (minor issues that don’t affect understanding). - **3**: Acceptable (noticeable issues but meaning is clear). - **2**: Poor (significant problems affecting comprehension). - **1**: Unacceptable (major errors, meaning unclear or wrong). ## Output Format Provide ONLY this JSON format (no additional text):
    "accuracy": <1-5>,
    "fluency": <1-5>,
    "completeness": <1-5>,
    "style_consistency": <1-5>,
    "comments": "Brief justification"
Now, let’s start this task: ## Input ### Original Text (English) {{ORIGINAL_TEXT}}
### Translation ({{LANGUAGE}}) {{TRANSLATED_TEXT}}
Model Arabic Chinese German English French Italian Korean Portuguese Russian Spanish
nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI
PosIR Multilingual Retrieval
gte-multilingual-base 34.08 0.52 51.52 0.26 47.96 0.44 61.59 0.19 49.59 0.41 47.63 0.45 37.54 0.48 48.94 0.42 44.89 0.46 49.95 0.42
bge-m3 34.08 0.41 47.93 0.34 46.59 0.29 50.79 0.30 44.07 0.33 42.50 0.34 38.19 0.35 44.27 0.36 40.60 0.36 43.16 0.35
Qwen3-Embedding-0.6B 42.11 0.45 57.49 0.30 53.79 0.37 65.10 0.22 55.33 0.34 54.49 0.36 45.68 0.38 55.68 0.35 50.20 0.40 56.43 0.33
inf-retriever-v1-1.5b 46.48 0.32 64.12 0.17 58.83 0.23 69.26 0.09 60.85 0.21 59.39 0.25 51.52 0.25 60.94 0.23 55.49 0.23 61.21 0.23
Qwen3-Embedding-4B 54.10 0.32 60.32 0.31 63.67 0.25 69.96 0.22 62.47 0.27 63.84 0.26 57.26 0.27 64.78 0.26 60.51 0.30 65.67 0.25
inf-retriever-v1 55.03 0.13 68.58 0.15 65.59 0.10 72.78 0.08 66.47 0.11 65.76 0.10 58.65 0.13 67.01 0.11 63.02 0.08 67.18 0.10
NV-Embed-v2 18.32 0.75 37.46 0.56 51.28 0.49 68.27 0.32 52.40 0.48 50.79 0.50 28.17 0.72 52.41 0.51 37.42 0.62 53.67 0.49
llama-embed-nemotron-8b 55.41 0.07 60.01 0.11 66.34 0.06 73.48 0.06 65.73 0.05 65.79 0.06 58.78 0.07 66.67 0.07 61.94 0.06 66.71 0.06
Qwen3-Embedding-8B 57.43 0.20 60.36 0.32 65.30 0.18 70.27 0.20 65.32 0.15 65.70 0.18 59.57 0.17 66.57 0.20 62.95 0.19 67.32 0.19
KaLM-Embedding-12B 43.74 0.11 50.82 0.11 53.33 0.08 64.22 0.06 52.40 0.06 52.15 0.07 43.80 0.10 51.71 0.07 51.73 0.08 54.79 0.06
Cross-lingual Retrieval
gte-multilingual-base 31.28 0.28 - - 50.75 0.26 - - 53.47 0.24 51.62 0.27 32.88 0.23 53.38 0.24 44.45 0.28 54.06 0.24
bge-m3 24.04 0.39 - - 43.16 0.35 - - 41.70 0.35 40.37 0.35 28.06 0.43 41.26 0.36 31.83 0.39 39.97 0.36
Qwen3-Embedding-0.6B 34.91 0.38 - - 53.97 0.29 - - 55.14 0.28 54.28 0.29 39.50 0.32 55.52 0.28 47.68 0.29 56.30 0.27
inf-retriever-v1-1.5b 35.68 0.23 - - 57.33 0.14 - - 60.08 0.12 58.19 0.13 42.74 0.18 60.14 0.13 50.34 0.17 60.39 0.13
Qwen3-Embedding-4B 52.20 0.32 - - 63.38 0.25 - - 62.88 0.27 63.49 0.26 55.74 0.28 64.09 0.27 59.50 0.28 64.80 0.26
inf-retriever-v1 52.28 0.17 - - 65.34 0.10 - - 66.77 0.10 65.89 0.10 56.62 0.14 66.94 0.10 61.82 0.12 67.28 0.10
NV-Embed-v2 20.60 0.34 - - 56.47 0.30 - - 58.42 0.30 57.20 0.30 20.29 0.39 58.75 0.30 37.48 0.34 58.52 0.30
llama-embed-nemotron-8b 52.73 0.08 - - 65.80 0.04 - - 66.18 0.06 66.21 0.05 56.35 0.08 66.74 0.06 60.47 0.05 66.94 0.05
Qwen3-Embedding-8B 53.97 0.26 - - 64.33 0.22 - - 64.47 0.22 64.20 0.23 57.28 0.25 64.92 0.23 60.81 0.24 65.47 0.22
KaLM-Embedding-12B 47.61 0.11 - - 57.65 0.07 - - 58.18 0.07 57.85 0.07 50.90 0.10 58.23 0.07 54.92 0.08 58.54 0.07
Q1(512) Multilingual Retrieval
gte-multilingual-base 49.55 0.33 61.52 0.22 61.47 0.21 73.24 0.07 64.12 0.20 61.56 0.22 54.06 0.28 62.82 0.19 59.86 0.24 64.62 0.18
bge-m3 46.29 0.40 62.37 0.34 59.89 0.21 66.00 0.23 58.53 0.28 56.63 0.29 52.32 0.32 58.36 0.28 53.41 0.31 57.82 0.29
Qwen3-Embedding-0.6B 50.25 0.35 69.10 0.19 62.12 0.21 77.34 0.10 63.80 0.21 62.89 0.18 54.66 0.23 64.43 0.21 59.04 0.29 65.70 0.17
inf-retriever-v1-1.5b 55.72 0.32 71.77 0.16 67.87 0.22 77.89 0.07 69.96 0.18 68.72 0.22 60.70 0.24 70.10 0.21 64.80 0.23 70.66 0.18
Qwen3-Embedding-4B 63.67 0.18 73.96 0.17 71.79 0.13 81.54 0.11 70.39 0.12 71.76 0.12 66.46 0.14 72.93 0.12 69.65 0.13 74.16 0.09
inf-retriever-v1 65.89 0.19 75.98 0.14 75.03 0.14 81.09 0.07 76.27 0.13 76.08 0.16 70.56 0.20 76.73 0.11 72.39 0.13 77.08 0.12
NV-Embed-v2 43.77 0.54 62.36 0.20 77.81 0.12 80.81 0.08 77.95 0.11 78.12 0.10 59.23 0.37 79.26 0.09 66.76 0.16 78.72 0.09
llama-embed-nemotron-8b 67.74 0.21 71.37 0.24 77.14 0.11 84.68 0.06 77.09 0.11 77.15 0.11 72.35 0.19 77.93 0.11 73.97 0.14 78.13 0.11
Qwen3-Embedding-8B 64.85 0.15 74.56 0.19 72.10 0.12 82.50 0.12 72.20 0.11 73.31 0.12 67.32 0.12 74.38 0.09 70.31 0.13 75.27 0.08
KaLM-Embedding-12B 68.66 0.18 66.89 0.12 76.11 0.12 79.16 0.07 75.71 0.10 75.72 0.11 70.58 0.16 75.76 0.09 74.48 0.10 77.04 0.07
Cross-lingual Retrieval
gte-multilingual-base 41.11 0.25 - - 60.45 0.14 - - 63.98 0.11 61.67 0.13 44.02 0.23 64.09 0.10 55.56 0.18 64.84 0.11
bge-m3 35.34 0.37 - - 54.93 0.31 - - 55.23 0.30 53.57 0.31 40.79 0.41 54.91 0.31 44.64 0.38 54.00 0.29
Qwen3-Embedding-0.6B 45.85 0.23 - - 65.33 0.11 - - 66.56 0.10 65.90 0.13 51.56 0.19 67.32 0.11 59.87 0.16 68.43 0.09
inf-retriever-v1-1.5b 43.60 0.30 - - 64.51 0.17 - - 68.19 0.13 66.04 0.16 51.07 0.26 68.50 0.14 59.11 0.18 68.84 0.12
Qwen3-Embedding-4B 65.11 0.14 - - 75.41 0.14 - - 74.92 0.12 75.53 0.13 69.10 0.13 76.11 0.13 72.29 0.12 76.99 0.12
inf-retriever-v1 57.93 0.20 - - 71.70 0.13 - - 73.67 0.11 72.51 0.11 63.78 0.16 73.74 0.09 68.28 0.14 74.21 0.10
NV-Embed-v2 20.78 0.33 - - 64.61 0.11 - - 68.83 0.13 65.27 0.12 23.73 0.26 69.30 0.10 41.04 0.16 67.86 0.09
llama-embed-nemotron-8b 65.40 0.15 - - 76.99 0.08 - - 77.66 0.09 77.59 0.09 70.22 0.15 78.29 0.07 73.22 0.12 78.62 0.07
Qwen3-Embedding-8B 68.63 0.12 - - 77.43 0.13 - - 77.57 0.12 77.53 0.12 72.25 0.13 78.27 0.12 74.80 0.12 78.73 0.12
KaLM-Embedding-12B 66.68 0.15 - - 74.57 0.11 - - 75.01 0.09 74.95 0.09 68.85 0.12 75.43 0.08 72.79 0.09 75.30 0.07
Q2(1024) Multilingual Retrieval
gte-multilingual-base 35.75 0.58 52.22 0.28 49.87 0.48 62.56 0.24 50.17 0.46 49.25 0.47 39.15 0.49 50.84 0.47 46.55 0.51 51.54 0.46
bge-m3 34.85 0.51 46.42 0.37 47.28 0.40 51.56 0.37 44.15 0.43 43.06 0.44 37.91 0.43 44.71 0.45 41.04 0.46 43.80 0.44
Qwen3-Embedding-0.6B 43.35 0.48 56.50 0.36 54.98 0.37 65.69 0.29 56.37 0.36 55.52 0.39 46.30 0.41 56.61 0.37 51.47 0.43 57.29 0.38
inf-retriever-v1-1.5b 46.65 0.39 62.60 0.22 58.51 0.30 69.11 0.17 60.21 0.27 58.98 0.29 50.97 0.30 60.83 0.30 55.46 0.28 60.68 0.26
Qwen3-Embedding-4B 55.19 0.36 59.42 0.32 64.72 0.28 70.65 0.24 63.50 0.29 64.69 0.28 57.69 0.30 65.58 0.29 61.23 0.31 66.40 0.28
inf-retriever-v1 55.56 0.25 67.03 0.22 65.75 0.21 72.93 0.16 66.37 0.21 66.08 0.19 58.47 0.25 67.28 0.22 63.62 0.19 67.24 0.21
NV-Embed-v2 11.80 0.84 41.32 0.61 54.51 0.47 72.56 0.11 57.53 0.43 55.16 0.53 24.85 0.77 56.90 0.47 37.44 0.67 59.14 0.43
llama-embed-nemotron-8b 55.91 0.22 59.46 0.20 66.67 0.16 73.56 0.14 65.87 0.18 65.97 0.16 58.45 0.19 66.80 0.18 62.00 0.17 66.88 0.16
Qwen3-Embedding-8B 58.28 0.26 59.76 0.31 65.66 0.23 71.31 0.21 65.61 0.19 65.96 0.21 59.36 0.21 66.67 0.21 63.13 0.25 67.58 0.21
KaLM-Embedding-12B 43.63 0.33 54.69 0.21 56.04 0.25 71.77 0.17 54.72 0.26 54.35 0.25 44.52 0.33 55.23 0.27 52.84 0.26 58.59 0.25
Cross-lingual Retrieval
gte-multilingual-base 31.88 0.41 - - 51.87 0.29 - - 54.34 0.28 52.38 0.30 33.13 0.39 54.10 0.27 45.32 0.34 54.80 0.30
bge-m3 23.61 0.56 - - 43.68 0.43 - - 41.91 0.42 40.63 0.46 27.17 0.51 41.57 0.42 31.54 0.53 40.26 0.45
Qwen3-Embedding-0.6B 34.76 0.47 - - 54.07 0.35 - - 55.18 0.35 54.22 0.33 39.07 0.41 55.64 0.34 47.54 0.40 56.29 0.33
inf-retriever-v1-1.5b 34.80 0.37 - - 57.08 0.26 - - 59.37 0.23 57.53 0.24 42.48 0.29 59.63 0.23 49.28 0.27 59.79 0.22
Qwen3-Embedding-4B 52.24 0.34 - - 63.89 0.29 - - 62.95 0.30 63.62 0.30 55.76 0.33 64.16 0.29 59.46 0.31 65.05 0.28
inf-retriever-v1 52.74 0.28 - - 65.53 0.19 - - 66.86 0.20 66.05 0.21 56.76 0.24 67.24 0.21 62.12 0.23 67.37 0.22
NV-Embed-v2 24.07 0.22 - - 60.85 0.13 - - 62.60 0.11 62.05 0.13 22.51 0.15 63.08 0.11 42.00 0.16 62.89 0.12
llama-embed-nemotron-8b 52.25 0.26 - - 65.85 0.18 - - 66.03 0.17 65.97 0.18 55.72 0.22 66.50 0.18 59.91 0.23 66.72 0.17
Qwen3-Embedding-8B 54.79 0.28 - - 65.10 0.25 - - 65.02 0.26 64.63 0.26 57.65 0.28 65.50 0.24 61.34 0.27 66.15 0.24
KaLM-Embedding-12B 56.99 0.27 - - 66.42 0.21 - - 66.66 0.21 66.38 0.21 59.67 0.24 67.06 0.21 63.86 0.23 66.89 0.21
Q3(1536) Multilingual Retrieval
gte-multilingual-base 24.71 0.74 45.26 0.29 40.32 0.58 54.46 0.32 41.32 0.58 40.12 0.62 27.32 0.70 41.18 0.59 36.08 0.62 41.30 0.60
bge-m3 27.38 0.61 39.08 0.40 38.73 0.43 42.27 0.44 35.67 0.48 34.03 0.51 30.06 0.50 36.39 0.50 33.23 0.49 34.44 0.52
Qwen3-Embedding-0.6B 37.40 0.57 49.86 0.39 49.06 0.47 57.28 0.39 50.53 0.46 49.69 0.48 40.35 0.53 51.00 0.47 44.60 0.52 51.30 0.47
inf-retriever-v1-1.5b 41.36 0.47 59.90 0.19 54.13 0.31 63.99 0.16 55.98 0.33 54.67 0.36 46.88 0.41 56.03 0.33 50.07 0.35 56.01 0.35
Qwen3-Embedding-4B 48.34 0.47 51.28 0.41 59.17 0.37 62.94 0.35 58.01 0.38 59.32 0.39 51.92 0.41 60.35 0.37 55.18 0.41 60.85 0.36
inf-retriever-v1 48.63 0.19 64.91 0.19 60.67 0.16 67.77 0.15 61.25 0.17 60.18 0.18 52.40 0.17 62.16 0.15 57.94 0.15 62.32 0.17
NV-Embed-v2 1.70 0.98 22.11 0.78 30.69 0.69 61.03 0.44 33.03 0.64 29.67 0.68 5.10 0.96 31.67 0.69 14.19 0.83 34.16 0.66
llama-embed-nemotron-8b 47.95 0.20 54.18 0.12 60.28 0.14 66.87 0.12 59.14 0.15 59.32 0.14 51.17 0.18 60.48 0.16 55.26 0.14 60.04 0.16
Qwen3-Embedding-8B 53.11 0.31 50.48 0.42 61.69 0.28 62.25 0.35 61.40 0.26 61.51 0.28 55.27 0.29 62.54 0.32 58.77 0.30 62.93 0.30
KaLM-Embedding-12B 26.45 0.30 37.66 0.28 36.71 0.22 50.08 0.23 35.00 0.27 34.90 0.30 24.21 0.36 32.93 0.26 35.65 0.25 37.52 0.30
Cross-lingual Retrieval
gte-multilingual-base 24.59 0.42 - - 45.21 0.34 - - 47.21 0.33 45.94 0.36 25.56 0.34 47.43 0.35 37.18 0.37 47.94 0.35
bge-m3 17.04 0.64 - - 36.77 0.46 - - 34.46 0.48 32.87 0.50 20.65 0.55 33.85 0.51 24.59 0.51 32.20 0.53
Qwen3-Embedding-0.6B 27.73 0.57 - - 47.01 0.46 - - 48.26 0.49 47.34 0.47 31.08 0.50 48.47 0.45 39.70 0.46 48.85 0.46
inf-retriever-v1-1.5b 30.59 0.27 - - 53.46 0.24 - - 55.78 0.22 53.70 0.23 37.07 0.29 55.22 0.25 45.03 0.20 55.60 0.24
Qwen3-Embedding-4B 44.18 0.47 - - 56.18 0.41 - - 55.86 0.45 56.46 0.43 47.26 0.39 57.26 0.43 51.86 0.44 57.72 0.42
inf-retriever-v1 49.09 0.23 - - 61.91 0.18 - - 63.02 0.18 62.29 0.16 51.99 0.19 63.36 0.18 58.40 0.19 63.62 0.18
NV-Embed-v2 19.80 0.61 - - 51.56 0.48 - - 51.97 0.47 52.47 0.47 17.34 0.64 52.57 0.46 34.65 0.54 53.02 0.49
llama-embed-nemotron-8b 44.91 0.19 - - 59.04 0.17 - - 59.26 0.13 59.46 0.14 47.49 0.14 59.94 0.18 52.96 0.17 59.91 0.15
Qwen3-Embedding-8B 43.82 0.47 - - 56.09 0.37 - - 56.22 0.39 55.92 0.39 47.01 0.43 56.41 0.42 51.55 0.40 57.03 0.39
KaLM-Embedding-12B 28.01 0.33 - - 40.94 0.26 - - 41.78 0.27 41.24 0.27 32.68 0.27 41.07 0.28 37.02 0.31 42.28 0.29
Table 10: Detailed evaluation results of 10 multilingual retrieval models on PosIR (Part 1 of 2). In both multilingual retrieval and cross-lingual retrieval (translated queries retrieving English documents) settings, the results for each language are weighted-averaged across 31 domains. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). “KaLM-Embedding-12B” denotes the “KaLM-Embedding-Gemma3-12B-2511” model. For Part 2, refer to Table 11.
Model Arabic Chinese German English French Italian Korean Portuguese Russian Spanish
nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI nDCG@10 PSI
Q4(2048) Multilingual Retrieval
gte-multilingual-base 18.41 0.78 39.21 0.40 32.10 0.67 48.46 0.38 35.11 0.62 31.76 0.67 20.19 0.74 32.57 0.64 28.78 0.69 33.50 0.63
bge-m3 22.82 0.48 34.22 0.42 32.62 0.42 34.30 0.40 30.98 0.40 28.76 0.46 25.69 0.44 30.57 0.47 28.17 0.42 29.08 0.46
Qwen3-Embedding-0.6B 34.18 0.57 43.84 0.44 44.13 0.51 50.39 0.39 45.78 0.47 45.08 0.49 37.03 0.52 45.55 0.48 41.43 0.53 45.79 0.49
inf-retriever-v1-1.5b 39.18 0.36 57.73 0.14 51.06 0.30 61.13 0.16 53.37 0.26 51.65 0.30 44.03 0.32 52.83 0.28 47.65 0.26 53.40 0.31
Qwen3-Embedding-4B 43.76 0.48 43.94 0.49 53.04 0.41 54.94 0.41 52.09 0.43 54.05 0.42 47.61 0.47 54.40 0.40 50.46 0.45 55.29 0.39
inf-retriever-v1 43.55 0.20 63.03 0.17 55.31 0.16 63.61 0.15 56.13 0.20 54.21 0.16 46.81 0.20 55.77 0.17 52.39 0.20 55.99 0.16
NV-Embed-v2 0.21 1.00 9.78 0.92 20.25 0.74 47.37 0.57 18.90 0.74 16.97 0.74 0.66 1.00 19.00 0.75 9.18 0.85 20.42 0.75
llama-embed-nemotron-8b 43.13 0.27 49.47 0.28 54.75 0.21 61.77 0.15 54.05 0.20 54.11 0.21 45.84 0.24 55.01 0.19 49.57 0.25 55.13 0.20
Qwen3-Embedding-8B 48.73 0.33 43.69 0.48 56.83 0.30 54.68 0.40 56.97 0.27 56.64 0.30 51.44 0.30 57.14 0.32 54.86 0.32 57.76 0.33
KaLM-Embedding-12B 20.88 0.40 30.88 0.29 27.99 0.29 41.33 0.17 27.51 0.30 26.83 0.36 18.55 0.37 24.85 0.35 28.67 0.34 29.00 0.30
Cross-lingual Retrieval
gte-multilingual-base 21.80 0.49 - - 39.50 0.41 - - 41.98 0.41 40.45 0.43 22.27 0.47 41.37 0.40 32.99 0.43 42.19 0.40
bge-m3 14.47 0.53 - - 31.06 0.44 - - 28.02 0.42 26.87 0.46 17.50 0.53 26.98 0.45 19.85 0.50 25.76 0.45
Qwen3-Embedding-0.6B 23.99 0.61 - - 41.08 0.48 - - 42.09 0.45 41.08 0.47 27.83 0.59 41.76 0.48 34.97 0.48 42.65 0.50
inf-retriever-v1-1.5b 30.48 0.24 - - 50.96 0.18 - - 53.05 0.14 51.61 0.13 36.60 0.27 52.93 0.17 43.84 0.20 53.06 0.19
Qwen3-Embedding-4B 37.33 0.50 - - 48.60 0.44 - - 48.43 0.45 48.86 0.44 40.74 0.52 49.31 0.46 44.62 0.48 49.84 0.44
inf-retriever-v1 46.25 0.18 - - 58.48 0.16 - - 59.01 0.16 58.27 0.16 49.89 0.19 58.80 0.13 54.42 0.16 59.36 0.14
NV-Embed-v2 16.81 0.66 - - 41.08 0.59 - - 40.98 0.60 40.78 0.59 14.19 0.70 40.48 0.59 27.57 0.64 41.57 0.59
llama-embed-nemotron-8b 41.60 0.19 - - 54.61 0.18 - - 54.99 0.17 55.18 0.17 44.35 0.20 55.45 0.17 48.75 0.19 55.58 0.20
Qwen3-Embedding-8B 37.03 0.47 - - 48.14 0.44 - - 48.58 0.42 48.04 0.45 40.78 0.48 48.71 0.45 44.48 0.45 49.36 0.45
KaLM-Embedding-12B 20.69 0.26 - - 32.23 0.22 - - 32.94 0.21 32.26 0.25 25.30 0.26 32.42 0.23 28.67 0.24 33.43 0.22
Table 11: Detailed evaluation results of 10 multilingual retrieval models on PosIR (Part 2 of 2). In both multilingual retrieval and cross-lingual retrieval (translated queries retrieving English documents) settings, the results for each language are weighted-averaged across 31 domains. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). “KaLM-Embedding-12B” denotes the “KaLM-Embedding-Gemma3-12B-2511” model.
Model Size Open-Sourced Model Link
Tokenizer
Qwen3 tokenizer Yang et al. (2025a) - https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507/blob/main/tokenizer.json
Embedding Model
gte-multilingual-base Zhang et al. (2024) 305M https://huggingface.co/Alibaba-NLP/gte-multilingual-base
bge-m3 Chen et al. (2024) 568M https://huggingface.co/BAAI/bge-m3
Qwen3-Embedding-0.6B Zhang et al. (2025) 595M https://huggingface.co/Qwen/Qwen3-Embedding-0.6B
inf-retriever-v1-1.5b Yang et al. (2025b) 2B https://huggingface.co/infly/inf-retriever-v1-1.5b
Qwen3-Embedding-4B Zhang et al. (2025) 4B https://huggingface.co/Qwen/Qwen3-Embedding-4B
inf-retriever-v1 Yang et al. (2025b) 7B https://huggingface.co/infly/inf-retriever-v1
QZhou-Embedding Yu et al. (2025) 7B https://huggingface.co/Kingsoft-LLM/QZhou-Embedding
NV-Embed-v2 Lee et al. (2025) 8B https://huggingface.co/nvidia/NV-Embed-v2
llama-embed-nemotron-8b Babakhin et al. (2025) 8B https://huggingface.co/nvidia/llama-embed-nemotron-8b
Qwen3-Embedding-8B Zhang et al. (2025) 8B https://huggingface.co/Qwen/Qwen3-Embedding-8B
KaLM-Embedding-Gemma3-12B-2511 Zhao et al. (2025a) 12B https://huggingface.co/tencent/KaLM-Embedding-Gemma3-12B-2511
Re-ranking Model
bge-reranker-v2-m3 Chen et al. (2024) 0.6B https://huggingface.co/BAAI/bge-reranker-v2-m3
Qwen3-Reranker-0.6B Zhang et al. (2025) 0.6B https://huggingface.co/Qwen/Qwen3-Reranker-0.6B
bge-reranker-v2-minicpm-layerwise Chen et al. (2024) 3B https://huggingface.co/BAAI/bge-reranker-v2-minicpm-layerwise
Qwen3-Reranker-4B Zhang et al. (2025) 4B https://huggingface.co/Qwen/Qwen3-Reranker-4B
Large Language Model
Hunyuan-MT-7B Zheng et al. (2025) 8B https://huggingface.co/tencent/Hunyuan-MT-7B
Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a) 31B https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507
DeepSeek-V3.1 DeepSeek-AI et al. (2025) 685B https://huggingface.co/deepseek-ai/DeepSeek-V3.1
GPT-4o OpenAI et al. (2024) - https://platform.openai.com/docs/models/gpt-4o
Online Service
Google Translate - https://translate.google.com
Table 12: Detailed information on all of the models appearing in our paper.
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 42,596 1511.2 166.0% 1458 24.4 193.6%
Chinese 55,969 1009.2 - 1488 11.8
German 42,690 1339.6 147.1% 1459 20.7 164.2%
English 42,708 910.6 - 1459 12.6
French 42,699 1342.6 147.4% 1459 22.0 174.6%
Italian 42,690 1360.8 149.4% 1459 21.5 170.0%
Korean 42,602 1543.8 169.5% 1456 26.2 207.4%
Portuguese 42,696 1264.2 138.8% 1459 19.8 156.5%
Russian 42,671 1503.1 165.1% 1456 24.4 193.1%
Spanish 42,672 1250.5 137.3% 1459 20.7 163.8%
Table 13: Statistics for domain: Accommodation Catering Hotel (Part 1 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 57,393 1520.6 165.7% 1339 26.0 193.7%
Chinese 63,398 933.5 - 1492 12.8
German 57,655 1394.7 152.0% 1340 22.9 170.7%
English 57,671 917.4 - 1340 13.4
French 57,661 1387.6 151.2% 1340 24.2 180.7%
Italian 57,658 1421.0 154.9% 1340 23.7 176.7%
Korean 57,545 1516.9 165.3% 1335 27.0 201.5%
Portuguese 57,649 1319.8 143.9% 1340 21.7 161.8%
Russian 57,643 1577.4 171.9% 1340 26.5 197.6%
Spanish 57,660 1312.5 143.1% 1340 22.7 169.4%
Table 14: Statistics for domain: Aerospace (Part 2 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 57,327 1550.9 166.8% 1613 26.0 195.8%
Chinese 64,363 930.0 - 1649 11.8
German 57,466 1423.2 153.1% 1615 23.1 173.9%
English 57,490 929.6 - 1615 13.3
French 57,469 1426.3 153.4% 1615 24.6 185.4%
Italian 57,477 1445.9 155.5% 1615 23.8 179.0%
Korean 57,362 1539.8 165.6% 1615 26.9 202.5%
Portuguese 57,474 1338.9 144.0% 1615 21.9 164.6%
Russian 57,438 1597.3 171.8% 1615 26.6 200.6%
Spanish 57,463 1332.1 143.3% 1608 23.0 172.9%
Table 15: Statistics for domain: Agriculture Forestry Animal Husbandry Fishery (Part 3 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 54,352 1460.8 160.0% 1079 26.2 207.6%
Chinese 57,698 1008.1 - 1419 12.2
German 54,918 1358.2 148.7% 1081 23.1 183.2%
English 54,969 913.1 - 1083 12.6
French 54,921 1331.2 145.8% 1083 24.2 191.9%
Italian 54,931 1356.8 148.6% 1081 23.6 187.0%
Korean 54,786 1423.8 155.9% 1081 25.8 204.9%
Portuguese 54,927 1255.0 137.4% 1083 21.0 166.8%
Russian 54,904 1435.1 157.2% 1083 25.4 201.6%
Spanish 54,912 1241.2 135.9% 1081 22.0 174.5%
Table 16: Statistics for domain: Artificial Intelligence Machine Learning (Part 4 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 55,866 1471.7 165.7% 1300 26.0 196.2%
Chinese 62,074 943.6 - 1329 12.3
German 55,929 1341.8 151.1% 1299 22.6 170.7%
English 55,948 888.1 - 1303 13.3
French 55,933 1312.9 147.8% 1303 24.0 180.6%
Italian 55,931 1367.2 154.0% 1303 23.6 178.0%
Korean 55,846 1452.9 163.6% 1301 27.3 205.7%
Portuguese 55,936 1257.3 141.6% 1303 21.5 162.2%
Russian 55,922 1462.4 164.7% 1303 25.6 192.6%
Spanish 55,932 1251.5 140.9% 1303 22.7 170.9%
Table 17: Statistics for domain: Automobile (Part 5 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 61,392 1546.3 167.9% 1406 27.7 200.9%
Chinese 63,133 930.7 - 1254 12.2
German 61,546 1405.8 152.7% 1411 23.8 172.9%
English 61,582 920.8 - 1411 13.8
French 61,562 1394.0 151.4% 1411 25.2 183.2%
Italian 61,562 1410.9 153.2% 1411 24.0 174.5%
Korean 61,465 1477.3 160.4% 1411 27.4 199.1%
Portuguese 61,562 1318.6 143.2% 1411 22.3 162.2%
Russian 61,545 1571.4 170.7% 1411 27.0 195.8%
Spanish 61,561 1308.2 142.1% 1411 23.3 169.1%
Table 18: Statistics for domain: Biomedicine (Part 6 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 61,435 1512.8 164.0% 1384 25.7 200.5%
Chinese 64,004 931.5 - 1657 11.6
German 61,852 1383.2 149.9% 1388 22.7 176.8%
English 61,885 922.6 - 1389 12.8
French 61,861 1352.0 146.5% 1387 23.5 183.5%
Italian 61,856 1385.5 150.2% 1388 22.8 177.6%
Korean 61,800 1500.8 162.7% 1387 26.6 207.0%
Portuguese 61,860 1285.5 139.3% 1386 20.9 162.6%
Russian 61,854 1480.0 160.4% 1389 24.6 191.8%
Spanish 61,847 1265.8 137.2% 1389 21.8 169.6%
Table 19: Statistics for domain: Computer Communication (Part 7 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 50,902 1515.4 154.0% 615 24.5 198.5%
Chinese 56,802 995.7 - 945 11.7
German 51,499 1394.2 141.7% 617 21.6 175.3%
English 51,612 983.9 - 624 12.3
French 51,511 1365.0 138.7% 624 22.0 178.6%
Italian 51,544 1383.9 140.7% 618 21.6 175.0%
Korean 51,385 1485.0 150.9% 617 24.5 198.8%
Portuguese 51,540 1284.6 130.6% 624 19.5 158.3%
Russian 51,524 1424.9 144.8% 615 22.8 184.8%
Spanish 51,496 1272.4 129.3% 623 20.5 166.4%
Table 20: Statistics for domain: Computer Programming Code (Part 8 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 60,224 1485.1 159.9% 1270 25.5 185.5%
Chinese 63,941 927.4 - 1280 12.3
German 60,374 1442.3 155.3% 1284 24.3 176.5%
English 60,390 928.9 - 1286 13.7
French 60,375 1423.4 153.2% 1286 25.3 184.0%
Italian 60,382 1445.6 155.6% 1286 24.5 178.3%
Korean 60,216 1536.2 165.4% 1282 27.8 202.0%
Portuguese 60,375 1329.8 143.2% 1286 22.3 162.6%
Russian 60,356 1604.8 172.8% 1286 27.9 203.3%
Spanish 60,365 1319.8 142.1% 1286 23.2 168.7%
Table 21: Statistics for domain: Current Affairs Government Administration (Part 9 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 59,471 1540.7 167.8% 1283 26.9 199.3%
Chinese 63,487 932.5 - 1535 12.5
German 59,557 1433.1 156.0% 1285 24.0 177.8%
English 59,580 918.4 - 1288 13.5
French 59,562 1436.6 156.4% 1288 25.6 189.3%
Italian 59,561 1469.3 160.0% 1288 25.0 184.8%
Korean 59,486 1526.2 166.2% 1288 27.5 204.0%
Portuguese 59,568 1346.8 146.7% 1288 22.6 167.4%
Russian 59,551 1612.0 175.5% 1288 27.9 206.6%
Spanish 59,566 1333.6 145.2% 1288 23.5 174.2%
Table 22: Statistics for domain: Electric Power Energy (Part 10 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 58,987 1530.5 164.5% 1679 30.6 227.9%
Chinese 63,899 932.2 - 1739 12.6
German 59,235 1330.3 143.0% 1687 21.4 159.4%
English 59,250 930.3 - 1687 13.4
French 59,236 1320.7 142.0% 1687 22.6 167.9%
Italian 59,236 1330.7 143.0% 1687 21.8 162.0%
Korean 59,079 1579.6 169.8% 1687 27.9 208.0%
Portuguese 59,235 1246.7 134.0% 1687 20.0 148.7%
Russian 59,205 1538.4 165.4% 1684 26.3 196.1%
Spanish 59,213 1241.6 133.5% 1687 21.1 157.3%
Table 23: Statistics for domain: Film Entertainment (Part 11 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 60,258 1518.8 163.5% 1315 26.6 193.5%
Chinese 64,184 931.8 - 1118 12.1
German 60,373 1443.0 155.4% 1315 24.3 176.4%
English 60,395 928.7 - 1315 13.8
French 60,375 1423.6 153.3% 1315 25.5 185.2%
Italian 60,383 1453.1 156.5% 1315 24.7 179.1%
Korean 60,287 1535.0 165.3% 1312 27.5 199.9%
Portuguese 60,382 1338.0 144.1% 1315 22.6 164.0%
Russian 60,358 1569.8 169.0% 1315 27.4 199.1%
Spanish 60,369 1320.5 142.2% $13$ 23.3 169.1%
Table 24: Statistics for domain: Finance Economics (Part 12 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 69,570 1512.8 163.5% $17$ 24.2 192.1%
Chinese 69,998 931.4 - 1218 12.3
German 69,918 1361.3 147.1% 1728 20.8 164.9%
English 69,999 925.1 - 1728 12.6
French 69,933 1348.3 145.7% 1728 22.1 175.4%
Italian 69,912 1368.2 147.9% 1728 21.5 170.7%
Korean 69,643 1537.2 166.2% $17$ 25.8 204.9%
Portuguese 69,939 1271.1 137.4% 1728 19.8 156.8%
Russian 69,838 1506.6 162.9% 1728 24.3 192.5%
Spanish 69,883 1258.1 136.0% 1728 20.7 164.2%
Table 25: Statistics for domain: Fineweb (Part 13 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 29,634 1573.6 166.6% 1294 25.7 196.6%
Chinese 46,710 966.3 - 1195 12.0
German 29,687 1455.5 154.1% 1295 23.4 178.7%
English 29,703 944.6 - 1296 13.1
French 29,700 1459.0 154.5% 1294 25.3 193.7%
Italian 29,696 1503.4 159.2% 1296 24.9 190.5%
Korean 29,644 1559.7 165.1% 1294 26.7 204.6%
Portuguese 29,696 1385.8 146.7% 1296 22.9 175.0%
Russian 29,684 1630.1 172.6% 1296 27.5 210.8%
Spanish 29,692 1360.0 144.0% 1296 23.3 178.2%
Table 26: Statistics for domain: Fire Safety Food Safety (Part 14 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 44,213 1549.6 163.5% 1276 24.7 197.1%
Chinese 44,119 1100.7 - 1499 12.1
German 44,341 1386.4 146.3% 1281 20.4 162.8%
English 44,373 947.7 - 1281 12.5
French 44,353 1361.2 143.6% 1281 21.5 171.4%
Italian 44,358 1397.4 147.4% 1281 20.8 166.0%
Korean 44,272 1591.5 167.9% 1278 26.3 210.1%
Portuguese 44,353 1292.4 136.4% 1281 19.3 153.6%
Russian 44,299 1490.8 157.3% 1275 23.0 183.6%
Spanish 44,316 1277.2 134.8% 1281 20.1 160.8%
Table 27: Statistics for domain: Game (Part 15 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 59,765 1514.9 163.7% 1287 26.6 195.7%
Chinese 64,276 931.6 - 1354 12.1
German 59,954 1453.8 157.1% 1283 24.6 180.4%
English 59,987 925.5 - 1293 13.6
French 59,962 1410.8 152.4% 1286 25.6 188.1%
Italian 59,965 1450.2 156.7% 1293 24.9 183.1%
Korean 59,773 1504.3 162.5% 1293 27.8 204.0%
Portuguese 59,959 1321.4 142.8% 1286 22.2 163.2%
Russian 59,950 1592.2 172.0% 1286 28.4 208.8%
Spanish 59,953 1312.7 141.8% 1286 23.2 170.4%
Table 28: Statistics for domain: Law Judiciary (Part 16 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 59,078 1499.8 160.2% 1579 24.3 180.3%
Chinese 64,410 935.4 - 1455 12.4
German 59,407 1352.2 144.5% 1586 20.8 154.4%
English 59,439 936.0 - 1586 13.5
French 59,415 1343.0 143.5% 1586 22.4 165.7%
Italian 59,410 1348.0 144.0% 1586 21.3 157.7%
Korean 59,189 1557.0 166.3% 1586 26.9 199.1%
Portuguese 59,408 1265.4 135.2% 1586 19.5 144.8%
Russian 59,380 1534.6 164.0% 1581 25.6 189.6%
Spanish 59,402 1257.8 134.4% 1586 20.7 153.5%
Table 29: Statistics for domain: Literature Emotion (Part 17 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 62,943 1398.9 149.4% 1032 26.1 186.5%
Chinese 50,306 1014.6 - 977 12.8
German 63,178 1307.3 139.6% 1041 24.0 171.6%
English 63,420 936.4 - 1041 14.0
French 63,214 1278.4 136.5% 1039 24.6 176.2%
Italian 63,196 1290.9 137.9% 1041 24.0 172.0%
Korean 63,049 1354.1 144.6% 1036 26.5 189.3%
Portuguese 63,201 1221.4 130.4% 1041 22.1 158.3%
Russian 63,188 1380.2 147.4% 1035 26.8 191.9%
Spanish 63,169 1216.4 129.9% 1041 23.1 165.4%
Table 30: Statistics for domain: Mathematics Statistics (Part 18 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 63,143 1554.0 167.8% 1385 27.5 204.1%
Chinese 64,519 929.2 - 1469 11.6
German 63,285 1426.6 154.0% 1385 24.4 180.7%
English 63,314 926.3 - 1385 13.5
French 63,297 1420.3 153.3% 1385 25.6 189.9%
Italian 63,292 1433.9 154.8% 1385 24.7 183.2%
Korean 63,197 1505.5 162.5% 1378 27.4 203.7%
Portuguese 63,297 1323.2 142.8% 1383 22.5 166.9%
Russian 63,271 1598.4 172.5% 1385 27.8 206.5%
Spanish 63,288 1320.7 142.6% 1385 23.7 175.7%
Table 31: Statistics for domain: Medicine Health Psychology Traditional Chinese Medicine (Part 19 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 47,403 1435.2 165.1% 1319 26.3 195.2%
Chinese 58,159 975.6 - 1474 12.6
German 47,500 1329.7 152.9% 1319 23.7 176.1%
English 47,521 869.5 - 1319 13.5
French 47,500 1325.8 152.5% 1319 24.8 183.9%
Italian 47,503 1351.8 155.5% 1319 24.0 178.4%
Korean 47,438 1448.9 166.6% 1319 33.9 251.5%
Portuguese 47,510 1254.8 144.3% 1319 22.2 164.6%
Russian 47,497 1495.8 172.0% 1319 27.5 204.3%
Spanish 47,497 1240.2 142.6% 1319 23.0 170.3%
Table 32: Statistics for domain: Mining (Part 20 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 53,784 1497.8 162.1% 1264 25.6 190.8%
Chinese 55,021 1024.2 - 1453 12.1
German 53,905 1376.1 148.9% 1270 22.6 168.3%
English 53,931 924.1 - 1271 13.4
French 53,918 1376.2 148.9% 1271 24.0 178.9%
Italian 53,922 1391.0 150.5% 1271 23.1 172.0%
Korean 53,785 1549.1 167.6% 1266 28.2 210.5%
Portuguese 53,918 1292.1 139.8% 1271 21.2 157.8%
Russian 53,897 1561.0 168.9% 1271 27.0 201.5%
Spanish 53,915 1278.5 138.3% 1271 22.0 163.8%
Table 33: Statistics for domain: News Media (Part 21 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 31,329 1571.8 165.4% 839 25.2 201.3%
Chinese 38,947 989.5 - 1044 11.8
German 31,360 1494.3 157.2% 839 22.8 182.5%
English 31,375 950.5 - 840 12.5
French 31,369 1478.1 155.5% 839 24.2 193.1%
Italian 31,366 1515.0 159.4% 839 23.6 188.7%
Korean 31,341 1566.9 164.8% 833 25.9 207.2%
Portuguese 31,367 1376.3 144.8% 839 21.1 168.4%
Russian 31,365 1635.4 172.1% 839 26.1 208.7%
Spanish 31,368 1359.9 143.1% 839 21.9 174.8%
Table 34: Statistics for domain: Other Information Services Information Security (Part 22 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 59,176 1472.4 163.3% 1609 25.8 196.9%
Chinese 63,160 928.7 - 1496 12.1
German 59,231 1370.3 151.9% 1615 22.9 174.5%
English 59,269 901.9 - 1615 13.1
French 59,248 1338.9 148.5% 1615 24.1 183.5%
Italian 59,250 1383.4 153.4% 1615 23.7 180.5%
Korean 59,170 1433.7 159.0% 1615 26.4 201.4%
Portuguese 59,253 1282.2 142.2% 1615 21.7 165.3%
Russian 59,235 1478.8 164.0% 1615 25.6 195.3%
Spanish 59,252 1268.2 140.6% 1615 22.7 172.7%
Table 35: Statistics for domain: Other Manufacturing (Part 23 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 57,032 1499.3 164.9% 1245 27.0 196.8%
Chinese 63,395 930.0 - 1605 12.7
German 57,116 1389.7 152.8% 1245 24.1 175.9%
English 57,150 909.4 - 1245 13.7
French 57,131 1393.0 153.2% 1245 25.7 187.0%
Italian 57,133 1421.4 156.3% 1245 24.9 181.3%
Korean 57,034 1487.5 163.6% 1240 27.7 201.9%
Portuguese 57,127 1317.5 144.9% 1245 23.3 169.5%
Russian 57,114 1559.4 171.5% 1243 27.8 202.9%
Spanish 57,116 1302.2 143.2% 1245 23.9 174.2%
Table 36: Statistics for domain: Petrochemical (Part 24 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 55,904 1529.9 166.1% 1618 25.2 191.3%
Chinese 63,618 931.3 - 1698 12.0
German 56,015 1403.1 152.3% 1618 22.7 172.1%
English 56,034 921.1 - 1618 13.2
French 56,024 1385.8 150.4% 1618 23.7 180.0%
Italian 56,018 1429.6 155.2% 1618 23.1 175.7%
Korean 55,928 1544.4 167.7% 1618 26.9 204.1%
Portuguese 56,020 1314.9 142.8% 1618 21.3 161.3%
Russian 55,999 1572.9 170.8% 1617 26.2 198.8%
Spanish 56,015 1307.3 141.9% 1618 22.2 168.7%
Table 37: Statistics for domain: Real Estate Construction (Part 25 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 55,871 1499.9 165.0% 1372 32.5 238.5%
Chinese 62,963 931.1 - 1429 12.3
German 56,023 1314.6 144.6% 1374 22.1 162.1%
English 56,048 908.8 - 1374 13.6
French 56,035 1323.3 145.6% 1374 23.8 174.6%
Italian 56,030 1336.0 147.0% 1374 22.9 167.9%
Korean 55,898 1542.0 169.7% 1368 34.4 252.4%
Portuguese 56,029 1246.1 137.1% 1374 21.2 155.8%
Russian 55,991 1519.3 167.2% 1374 33.4 244.9%
Spanish 56,024 1243.1 136.8% 1374 22.4 164.2%
Table 38: Statistics for domain: Sports (Part 26 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 61,343 1531.2 164.5% 1669 25.5 190.1%
Chinese 64,739 932.0 - 1751 11.4
German 61,544 1410.7 151.5% 1672 22.7 169.2%
English 61,577 931.0 - 1672 13.4
French 61,554 1408.9 151.3% 1672 24.3 180.9%
Italian 61,549 1432.4 153.9% 1672 23.2 172.7%
Korean 61,466 1549.0 166.4% 1672 27.1 202.3%
Portuguese 61,561 1318.0 141.6% 1672 21.1 157.2%
Russian 61,525 1590.1 170.8% 1672 26.4 197.0%
Spanish 61,551 1308.9 140.6% 1672 22.1 164.4%
Table 39: Statistics for domain: Subject Education Education (Part 27 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 61,031 1510.4 162.8% 1398 26.8 196.6%
Chinese 63,950 930.5 - 1594 12.0
German 61,220 1396.1 150.5% 1401 23.7 173.7%
English 61,254 927.7 - 1401 13.6
French 61,229 1381.2 148.9% 1401 25.0 183.3%
Italian 61,233 1405.4 151.5% 1401 24.1 176.9%
Korean 61,141 1490.1 160.6% 1401 27.4 201.0%
Portuguese 61,220 1311.7 141.4% 1401 22.2 163.0%
Russian 61,201 1550.6 167.1% 1399 27.2 199.7%
Spanish 61,218 1299.2 140.0% 1401 23.3 170.6%
Table 40: Statistics for domain: Technology Scientific Research (Part 28 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 53,946 1499.7 163.8% 1504 23.6 182.2%
Chinese 62,844 935.1 - 1456 11.7
German 54,127 1338.0 146.2% 1504 20.5 157.9%
English 54,144 915.4 - 1504 13.0
French 54,127 1346.8 147.1% 1504 22.0 169.7%
Italian 54,124 1363.1 148.9% 1504 21.1 163.2%
Korean 53,981 1524.5 166.5% 1487 25.2 194.7%
Portuguese 54,124 1279.5 139.8% 1504 19.7 152.2%
Russian 54,037 1538.5 168.1% 1503 24.5 188.9%
Spanish 54,099 1277.9 139.6% 1504 20.8 160.2%
Table 41: Statistics for domain: Tourism Geography (Part 29 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 53,730 1516.6 166.7% 1418 26.1 195.6%
Chinese 63,115 931.9 - 1370 12.2
German 53,812 1401.2 154.0% 1418 23.4 174.8%
English 53,829 909.9 - 1418 13.4
French 53,820 1383.5 152.1% 1418 24.4 182.9%
Italian 53,821 1435.5 157.8% 1418 24.4 182.4%
Korean 53,745 1513.9 166.4% 1415 27.3 204.5%
Portuguese 53,821 1317.5 144.8% 1418 22.0 164.8%
Russian 53,801 1580.0 173.6% 1418 27.3 204.2%
Spanish 53,818 1311.4 144.1% 1418 23.1 172.7%
Table 42: Statistics for domain: Transportation (Part 30 of 31).
Language #Corpus
Avg Token
of Corpus
Ratio
(Rel. English)
#Queries
Avg Token
of Queries
Ratio
(Rel. English)
Arabic 53,749 1514.7 166.9% 1328 26.3 195.4%
Chinese 63,027 929.0 - 1489 12.2
German 53,839 1403.8 154.7% 1328 23.8 176.6%
English 53,854 907.5 - 1328 13.5
French 53,838 1415.5 156.0% 1328 25.4 188.7%
Italian 53,842 1439.7 158.6% 1328 24.8 184.4%
Korean 53,746 1487.7 163.9% 1325 26.7 198.5%
Portuguese 53,841 1325.7 146.1% 1328 22.5 167.0%
Russian 53,813 1579.1 174.0% 1328 27.4 203.2%
Spanish 53,843 1323.7 145.9% 1328 23.5 174.7%
Table 43: Statistics for domain: Water Resources Ocean (Part 31 of 31).